### v1

# Communication-Efficient Wireless Federated Fine-Tuning for Large-Scale AI Models

Abstract
Transformer-based large language models (LLMs) have achieved remarkable success across various tasks. Yet, fine-tuning such massive models in federated learning (FL) settings poses significant challenges due to resource constraints and communication overhead. Low-Rank Adaptation (LoRA) addresses these issues by training compact, low-rank matrices instead of fully fine-tuning large models. This paper introduces a wireless federated LoRA fine-tuning framework that optimizes both learning performance and communication efficiency. We provide a novel convergence analysis, revealing how LoRA rank and covariance effects influence FL training dynamics. Leveraging these insights, we propose Sparsified Orthogonal Fine-Tuning (SOFT), an adaptive sparsification method that streamlines parameter updates without expensive matrix multiplications and singular value decomposition (SVD) operations. Additionally, we present a Two Stage Federated Algorithm (TSFA) that pre-determines LoRA rank offline and dynamically adjusts other parameters online, ensuring efficient training under latency constraints. Our framework thus enables scalable, resource-efficient deployment of large models in wireless FL scenarios.
Index Terms:
Parameter-efficient fine-tuning, federated learning, sparsification, Lyapunov optimization.
I Introduction
Transformer-based large language models (LLMs) like ChatGPT [1] have demonstrated exceptional performance across a multitude of tasks, including applications in Natural Language Processing (NLP), Computer Vision (CV), Healthcare, and Biomedical research. Despite their impressive capabilities, training these models from scratch is highly resource-intensive, demanding substantial time and computational power. Consequently, it has become common practice to fine-tune pre-trained models to cater to specific domains, maximizing efficiency while leveraging existing architectures.
As a promising solution , parameter-efficient fine-tuning (PEFT) methods have been developed [2, 3]. PEFT techniques focus on adjusting a smaller subset of model parameters, thereby reducing computational load and memory requirements while achieving performance comparable to that of fully fine-tuned models. Among these methods, Low-Rank Adaptation (LoRA) [3] stands out as a particularly effective approach. LoRA minimizes the number of trainable parameters by introducing trainable low-rank matrices integrated with the existing pre-trained model weights, which remain frozen during fine-tuning. This technique allows for efficient fine-tuning by updating only these additional low-rank matrices, significantly reducing computational and memory demands without substantially compromising performance. On many tasks, LoRA has been shown to achieve comparable or even superior performance compared to fully fine-tuning. As a result, LoRA facilitates the deployment of advanced LLMs in environments with limited resources, enhancing the practicality and accessibility of sophisticated capabilities.
In many practical scenarios, the data necessary for fine-tuning is distributed across multiple clients, and centralizing this data raises significant privacy and security concerns. To address these challenges, Federated Learning (FL) has emerged as a promising paradigm for collaborative model training without the need for centralized data aggregation [4, 5, 6, 7]. FL enables multiple clients to jointly train a global model by sharing model updates. Each client trains the model locally using its own data and periodically communicates the model parameters or gradients to a central server, which aggregates them to update the global model. In practical wireless FL systems, however, communication efficiency becomes critical due to limited bandwidth and fluctuating channel conditions [8]. To mitigate these challenges, sparsification techniques are commonly employed in FL by transmitting only the most significant model updates, typically by selecting the top-$k$ norm elements [9].
Several recent studies have attempted to combine LoRA with FL to harness the benefits of both methodologies [10, 11, 12]. For instance, the authors in [10] pioneered the integration of LoRA-based local updates with the Federated Averaging (FedAvg) algorithm [4] for model aggregation. Building upon this, [11] proposed a method for handling heterogeneous LoRA ranks among clients, suggesting that the server zero-pads the missing components corresponding to absent rank indices in clients with smaller ranks so that all matrices match the largest rank prior to aggregation. Meanwhile, [12] introduced Federated Freeze A LoRA (FFA-LoRA), a method that addresses data heterogeneity and provides privacy guarantees through differential privacy. FFA-LoRA uses only one of the two LoRA matrices during training and employs the Gaussian mechanism to analyze differential privacy. Finally, the authors in [13] investigated the impact of data heterogeneity on learning efficiency and introduced a data-driven initialization method to address this challenge.
Despite these advancements, they share a certain limitation. All the aforementioned works rely on a determined rank and do not provide a systematic method for selecting an optimal rank that balances performance and resource constraints. Since the performance of LoRA is highly sensitive to the choice of rank $r$ [3], this oversight leads to suboptimal model performance or unnecessary resource consumption. A rank that is too low may fail to capture essential features, leading to suboptimal model performance, while a higher rank makes the clients communicate larger model sizes with the server.
Moreover, a separate challenge arises when integrating sparsification methods in LoRA-based FL. In LoRA, model updates are factored into two low-rank matrices. Naively applying top-$k$ selection to each matrix independently disregards the fact that small entries in one matrix can have a disproportionately larger influence than other entries when multiplied by large entries in the other. To accurately capture the most influential updates, one would need to compute the full product of the two matrices, then apply a singular value decomposition (SVD) and transmit only the components corresponding to the largest singular values. However, frequent SVD operations are computationally prohibitive, especially for resource-constrained clients, making both naive top-$k$ and SVD-based approaches infeasible in LoRA-based FL. This underscores the need for novel sparsification methods that can select significant information while keeping computational costs feasible.
To address these challenges, we propose a practical wireless federated LoRA fine-tuning framework tailored for resource-constrained clients. Our approach dynamically optimizes the LoRA rank to balance learning performance with communication efficiency, and introduces a novel LoRA sparsification strategy, specifically not requiring full matrix multiplications and SVD operations. We specifically propose Sparsified Orthogonal Fine-Tuning (SOFT), which leverages the orthogonality constraints on LoRA matrices. In this scheme, each rank vector—comprising the corresponding column vectors in the left LoRA matrix and row vectors in the right LoRA matrix—is orthogonal. As a result, the first LoRA matrix effectively functions similarly to the left singular matrix in an SVD, while the second corresponds to the right singular matrix. To determine the importance of the $i$-th rank vector, we compute the product of the norm of its column in the left matrix and the norm of its row in the right matrix, using this result as a proxy for the singular value. By enforcing these orthogonality constraints, SOFT efficiently select the most significant updates while maintaining low computational overhead.
On the other hand, to determine LoRA rank, we perform a convergence analysis of LoRA within the FL setting. By incorporating LoRA-rank effects into our analysis, we derive theoretical insights into how this parameter affects the model performance and convergence speed. Based on this analysis, we first divide the optimization parameters into two groups: parameters that must be determined before the training begins and parameters that can be adjusted dynamically during the training. Specifically, since the LoRA rank represents the model structure used in FL training which must be established prior to the training, we determine it through an offline stage. After training commences, we adaptively adjust the sparsification ratios and bandwidth allocation strategy in response to real-time network conditions and resource availability in an online stage. To manage long-term latency constraints and ensure stable convergence, we employ Lyapunov optimization techniques.
Contributions.  The key contributions of this article are summarized as follows:
  • We introduce a new LoRA sparsification technique, SOFT, that avoids expensive computational operations by enforcing orthogonality in the LoRA matrices. Through an orthogonal regularization term in the loss function, SOFT ensures that these rank matrices behave similarly to left and right singular matrices in an SVD.
  • We conduct theoretical convergence analysis that explicitly incorporates the LoRA rank into FL for the first time. This analysis offers insights into how rank selection impacts model accuracy, sparsification error, and convergence speed.
  • We propose a Two Stage Federated Algorithm (TSFA) framework that first determines the LoRA rank under approximate channel conditions before the training begins, and then adaptively adjusts the sparsification ratio and bandwidth allocation at each iteration using Lyapunov optimization.
  • We rigorously validate the effectiveness of the proposed framework through extensive simulations and experiments on benchmark datasets. The results demonstrate that our approach achieves comparable or superior performance to the existing methods while significantly reducing communication overhead, thereby substantiating its practicality and efficiency in real-world FL applications.
Organization.  The remainder of this article is organized as follows. In Section II, we describe the system model of the proposed framework, including the LoRA mechanism, FL algorithm, and communication model. Section III introduces the sparsification method. In Section IV, we present the offline and online optimization approach for parameter selection. Section V provides numerical results demonstrating the effectiveness of our method. Finally, Section VI concludes the paper. Table I provides a summary of the key notations used.
II System Model
TABLE I: Notations and Descriptions
Notation Description
$\theta_{P}$ Pre-trained model weight matrix in $\mathbb{R}^{d\times\ell}$
$p_{k}$ Data proportion for client $k$, $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$
$\mathcal{D}_{k}$ Local dataset of client $k$
$B$ Total available bandwidth
$b_{k}^{t}$ Bandwidth allocation ratio for client $k$ at iteration $t$
$h_{k}^{t}$ Channel gain for client $k$ at iteration $t$
$D^{t}$ Overall transmission delay at iteration $t$
$\bar{D}$ Upper bound of $D^{t}$
$\mathcal{K}^{t}$ Set of scheduled clients at iteration $t$
$m_{k}^{t}$ Error feedback memory matrix at client $k$
$\tilde{m}_{k}^{t}$ Concatenated error feedback vector
$O^{t}$ Sparsification ratio at iteration $t$
$W$ Upper bound on the singular values of LoRA
$\phi$ Constant used in the covariance bound assumption
$S$ Smoothness constant of the loss functions
$G_{k}$ Upper bound on the gradient norm at client $k$
$G$ Maximum gradient bound, $G=\max_{k}G_{k}$
$Q^{t}$ Virtual queue at iteration $t$
$D_{th}$ Average transmission delay threshold
II-A Low-rank Adaptation
Suppose we are given a pre-trained model $\theta_{P}\in\mathbb{R}^{d\times\ell}$. We aim to adapt this pre-trained model through fine-tuning for use in downstream tasks. In traditional fine-tuning, the model is initialized with the pre-trained weights $\theta_{P}$ and updated via gradient descent to obtain $\theta=\theta_{P}+\Delta\theta$, where $\Delta\theta$ represents the updates to the weights over all $d\times\ell$ parameters.
Rather than updating $\Delta\theta$ directly, LoRA seeks to represent the weight updates as a low-rank decomposition $\Delta\theta=\theta_{B}\theta_{A}$, where $\theta_{B}\in\mathbb{R}^{d\times r}$, $\theta_{A}\in\mathbb{R}^{r\times\ell}$, and $r\ll\min(d,\ell)$. In this way, the output can be represented as $\boldsymbol{y}=\boldsymbol{x}\theta_{P}+\frac{\alpha}{r}\boldsymbol{x}\Delta\theta=\boldsymbol{x}\theta_{P}+\frac{\alpha}{r}\boldsymbol{x}\theta_{B}\theta_{A},$ where $\boldsymbol{y}\in\mathbb{R}^{\ell\times 1}$, $\boldsymbol{x}\in\mathbb{R}^{d\times 1}$, and $\frac{\alpha}{r}$ is a scaling parameter. In LoRA, the pre-trained weights $\theta_{P}$ are kept fixed during fine-tuning, and only the low-rank matrices $\theta_{A}$ and $\theta_{B}$ are updated. This significantly reduces the number of trainable parameters by a factor of $\mathcal{O}\left(\frac{r}{\min(d,\ell)}\right)$ compared to full fine-tuning.
The choice of the rank $r$ determines the fine-tuning performance, as it controls the level of approximation. A higher rank allows the model to capture more information from the original weight updates, potentially improving performance. However, it is observed that over-parameterized models often reside in a low intrinsic dimension [14, 3], meaning that there exists a maximum rank $r$ which can suffice to achieve performance comparable to fully fine-tuning.
To ensure that the weights start from the pre-trained model, $\theta_{B}$ is initialized to zero, while $\theta_{A}$ is initialized with random Gaussian values. The scaling parameter $\alpha$ controls the update magnitude, aiding in stable training. By updating only the low-rank matrices, LoRA achieves significant memory and computational efficiency, making it practical for fine-tuning large models on resource-constrained clients.
II-B Federated LoRA Fine-Tuning
Figure 1: System model of Federated LoRA Fine-Tuning.
Consider a network composed of a central parameter server and total $N$ local clients with single-antenna to fine-tune a global model $\theta\in\mathbb{R}^{d\times l}$ based on their local datasets using LoRA module with rank $r$. In the context of applying LoRA to FL, the pre-trained model remains fixed on both the server and the clients, so there is no need to transmit its parameters. Instead, only the LoRA $\theta_{B}$ and $\theta_{A}$ are updated locally and exchanged during each communication round. This approach significantly reduces communication overhead compared to fully fine-tuning methods that require transmitting the entire model at every round.
Each client $k\in\mathcal{N}$ owns a local dataset $\mathcal{D}_{k}=\{(x_{i},y_{i})\}_{i=1}^{D_{k}}$ with $D_{k}=|\mathcal{D}_{k}|$ data pairs such that $\mathcal{D}=\cup_{k\in\mathcal{N}}\mathcal{D}_{k}$. The local loss function of client $k$ is $F_{k}(\theta_{B},\theta_{A})=\frac{1}{D_{k}}\sum_{x_{i}\in\mathcal{D}_{k}}f(\theta_{B},\theta_{A};x_{i},y_{i}),$ where $f(\theta_{B},\theta_{A})$ represents the sample-wise loss function. The global loss function is $F(\theta_{B},\theta_{A})=\sum_{k\in\mathcal{K}^{t}}p_{k}F_{k}(\theta_{B},\theta_{A}),$ where $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$. To achieve a global model $\theta^{*}=\theta_{P}+\frac{\alpha}{r}\theta_{B}^{*}\theta_{A}^{*}$, clients periodically send updates to the server, assuming no direct interaction among the clients. The optimal global model $\theta^{*}$ for LoRA is identified through the following minimization process $\theta^{*}\!=\!\underset{\theta_{B}\in\mathbb{R}^{d\times r},\theta_{A}\in\mathbb{R}^{r\times\ell}}{\arg\min}F(\theta_{B},\theta_{A}).$ Within the wireless federated LoRA fine-tuning framework, the objective function $F(\theta_{B},\theta_{A})$ is progressively minimized through a series of local and global updates.
The aggregation process closely mirrors that of FedAvg, with the key difference being that only the LoRA modules are transmitted and aggregated. Specifically, the following steps are executed in the $t$-th round for $t=1,2,\dots,T$:
  1. 1.
    Global Matrices Broadcasting: The server distributes the $t$-th round global matrices $\theta_{B}^{t}$ and $\theta_{A}^{t}$ to the set of scheduled clients $\mathcal{K}^{t}$ by broadcasting. Since the server has enough power to transmit the global matrices correctly, we assume the downlink transmission is error-free.
  2. 2.
    Local Fine-Tuning: Each client $k\in\mathcal{K}^{t}$ initialize the global matrices as an initial point of local matrices, i.e. $\theta_{B,k}^{t}=\theta_{B}^{t},\theta_{A,k}^{t}=\theta_{A}^{t}$. Each client performs local training through gradient descent or ADAM resulting in an updated local models $\theta_{B,k}^{t+1}$ and $\theta_{A,k}^{t+1}$.
  3. 3.
    Matrices Aggregation: Client $k\in\mathcal{K}^{t}$ returns the updated local matrices $\theta_{B,k}^{t+1}$ and $\theta_{A,k}^{t+1}$ to the server. The server aggregates and averages the updates to form the $t+1$-th round global matrices as follows:
    $\theta_{B}^{t+1}=\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{B,k}^{t+1},\quad\theta_{A}^{t+1}=\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{A,k}^{t+1},$ (1)
where $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$ denotes the proportion of the dataset associated with client $k$, i.e. $\sum_{k\in\mathcal{N}}p_{k}=1$. After a total of $T$ rounds of this process, the training completes, yielding an optimally trained global model for LoRA. Although a client may intend to send the complete matrix product $\theta_{B,k}\theta_{A,k}$, to avoid the computational burden of matrix multiplication at the resource-constrained clients, it instead sends $\theta_{B,k}$ and $\theta_{A,k}$. This is because transmitting the full product would involve sending a matrix of size $d\times\ell$, which is generally much larger than $\theta_{B,k}$ and $\theta_{A,k}$ due to $r\ll\min\{d,\ell\}$. Therefore sending the two low-rank matrices not only minimizes local computation but also reduces the overall communication latency and overhead, as the total number of transmitted parameters is significantly low. Fig. 1 provides an overall framework.
II-C Communication Model
We consider a wireless multiple access system, specifically a frequency division multiple access (FDMA) scheme for local model transmission, with a total available bandwidth of $B$. The channel gain between the server and client is modeled as $h_{k}^{t}=s_{k}^{t}q_{k}^{-\gamma}$, where $s_{k}^{t}\sim\mathcal{N}(0,1)$ represents the small-scale channel fading coefficient, and $q_{k}^{-\gamma}$ denotes the distance-dependent path loss with exponent $\gamma$. Additionally, $b_{k}^{t}\in[0,1]$ represents the bandwidth allocation ratio for client $k\in\mathcal{N}$ where $\sum_{k=1}^{N}b_{k}^{t}=1$. The variable $a_{k}^{t}\in\{0,1\}$ indicates whether client $k$ is selected in iteration $t$, where $\sum_{k=1}^{N}a_{k}^{t}=K^{t}$ and $a_{k}^{t}=0$ implies no bandwidth is allocated to client $k$, resulting in $b_{k}^{t}=0$. The transmission rate for client $k$ can be expressed as $r_{k}^{t}=b_{k}^{t}B\log_{2}\left(1+\frac{|h_{k}^{t}|^{2}}{\sigma^{2}}\right)$, where $\sigma^{2}$ represents the variance of the additive white Gaussian noise (AWGN). During the uplink phase, we assume that the amount of information per model parameter is denoted by $v$. Thus, the communication delay for client $k$ at iteration $t$ can be written as $D_{k}^{t}=\frac{a_{k}^{t}vr(d+\ell)}{r_{k}^{t}}=\frac{a_{k}^{t}vr(d+\ell)}{b_{k}^{t}B\log_{2}\left(1+\frac{|h_{k}^{t}|^{2}}{\sigma^{2}}\right)}.$ Note that if $a_{k}^{t}=0$, then $D_{k}^{t}=0$. Since the overall transmission delay at iteration $t$ is determined by the slowest client, it is given by $D^{t}=\max_{k\in\mathcal{K}^{t}}D_{k}^{t}.$
III Sparsified Orthogonal Fine-Tuning
III-A Loss Function Design
A fundamental component of our sparsification method is the design of a loss function that enforces orthogonal properties on the matrices $\theta_{B,k}$ and $\theta_{A,k}$. This design is inspired by the orthogonality inherent in the singular matrices of SVD. Specifically, we aim for $\theta_{B,k}^{\textsf{T}}\theta_{B,k}$ and $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ to approximate diagonal matrices, analogous to the roles of $U$ and $V$ in SVD. To accomplish this, each client $k$ constructs the following loss function:
$L_{k}=L_{task,k}+\zeta(\|\theta_{B,k}^{\textsf{T}}\theta_{B,k}-diag(\theta_{B,k}^{\textsf{T}}\theta_{B,k})\|_{F}^{2}$
$+\|\theta_{A,k}\theta_{A,k}^{\textsf{T}}-diag(\theta_{A,k}\theta_{A,k}^{\textsf{T}})\|_{F}^{2}),$ (2)
where $L_{task,k}$ denotes the task-specific loss for client $k$, $\zeta$ is a scaling parameter, and $\text{diag}(C)$ denotes a diagonal matrix with the diagonal elements of $C$.
Fig. 2 illustrates the impact of our proposed loss function on promoting orthogonality within $\theta_{B,k}$ and $\theta_{A,k}$. After training with this loss function, the product $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ becomes approximately diagonal, demonstrating the successful enforcement of orthogonal properties. This property allows us to compute the Frobenius norm by aggregating the $\ell_{2}$-norms of the individual vectors across the $r$ components, as expressed mathematically $\|\theta_{B,k}\theta_{A,k}\|_{F}^{2}=\sum_{i=1}^{r}\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2},$ where $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$ are the $i$-th column and row vectors, respectively, whose orthogonality ensures that cross terms vanish, simplifying the calculation to sum of vector norms.
Figure 2: Visualization of $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ of Vit-Base blocks.6.mlp.fc1 layer: without (a) and with (b) the proposed loss function.
III-B Sparsification Method
At the beginning of the iteration, each client $k$ performs local training using the proposed loss function (III-A), resulting in $\theta_{B,k}$ and $\theta_{A,k}$. The client then applies sparsification to updated matrices. Specifically, the client sets the magnitude of all elements in $\theta_{B,k}$ and $\theta_{A,k}$ to zero, except for the selected $r(d+\ell)O_{k}$ elements, where $O_{k}\in[0,1]$ is the sparsification ratio of client $k$.
Since $\theta_{B,k}$ and $\theta_{A,k}$ exhibit orthogonal properties, we can compute singular values by calculating the norms of the matrices $\theta_{B,k}$ and $\theta_{A,k}$. For instance, the $i$-th singular value, where $i=\{1,...,r\}$, can be obatained as $\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2}$. Because the magnitude of singular values represents the importance of their corresponding singular vectors, we implement sparsification based on the magnitude of these singular values. For each singular vector, we define $o_{k,i}$ as the number of non-zero elements retained in both $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$. The value of $o_{k,i}$ is determined by distributing the total number of non-sparse elements proportionally to the magnitude of the singular values. Specifically, $o_{k,i}$ is calculated as follows:
$o_{k,i}=\frac{O_{k}r(d+\ell)\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2}}{\sum_{i=1}^{r}\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2}}.$ (3)
Finally, based on (3), client $k$ selects top-$o_{k,i}$ elements in $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$. Although the proposed sparsification method also requires calculating the $l$-2 norm $r$ times in each iteration, this computation has a complexity of $\mathcal{O}(r(d+\ell))$, negligible compared to the original complexity caused by full matrix multiplication and SVD operation, i.e. $\mathcal{O}(dr\ell+\min(d,\ell)\cdot d\ell)$.
Algorithm 1 Sparsified Orthogonal Fine-Tuning
1:Input: The global model $\theta_{B}$ and $\theta_{A}$, sparsification ratio $O_{k}$.
2:for $k\in\mathcal{K}^{t}$ $\triangleright$ Local update in parallel
3:    Download $\theta_{B_{k}}^{t,0}=\theta_{B}$ and $\theta_{A_{k}}^{t,0}=\theta_{A}$ from the PS
4:    for $e=0,\dots,E-1$ do
5:       Locally iterate using (III-A)
6:    end for
7:    Incorporate error feedback:
8:    $m_{B,k}^{t}+\Delta\theta_{B,k}^{t},m_{A,k}^{t}+\Delta\theta_{A,k}^{t}$
9:    Compute sparified updates $\theta_{B,k}$ and $\theta_{A,k}$ with rate $O_{k}$
10:    Transmit $\theta_{B,k}^{t}$ and $\theta_{A,k}^{t}$ to the server
11:end for
III-C Error Feedback with SOFT
While the proposed sparsification method effectively reduces communication costs by transmitting only the most significant elements of $\theta_{B,k}$ and $\theta_{A,k}$, it inevitably introduces sparsification errors. To mitigate the impact of these errors on model convergence and performance, we incorporate an error feedback mechanism into SOFT. The error feedback mechanism works by accumulating the sparsification errors at each client and incorporating them into the updates [9]. Specifically, each client maintains local error memory matrices, denoted as $m_{B,k}$ and $m_{A,k}$, which store the cumulative sparsification errors for $\theta_{B,k}$ and $\theta_{A,k}$, respectively. The sparified update can be expressed as $\theta_{B,k}^{t}=\mathcal{S}(m_{B,k}^{t}+\Delta\theta_{B,k}^{t}),\theta_{A,k}^{t}=\mathcal{S}(m_{A,k}^{t}+\Delta\theta_{A,k}^{t}),$ where $\mathcal{S}$ operates the sparsification algorithm, $\Delta\theta_{B,k}^{t}$ and $\Delta\theta_{A,k}^{t}$ represent obtained model of client $k$ at time $t$ using the loss function (III-A). Subsequently, the error memory is updated to accumulate the residuals resulting from sparsification:
$m_{B,k}^{t+1}=m_{B,k}^{t}+\Delta\theta_{B,k}^{t}-\theta_{B,k}^{t},$ (4)
$m_{A,k}^{t+1}=m_{A,k}^{t}+\Delta\theta_{A,k}^{t}-\theta_{A,k}^{t}.$ (5)
For clarity and brevity, we define the concatenated representations as follows:
$\tilde{m}_{k}^{t}\!=\!\begin{bmatrix}{m_{B,k}^{t\textsf{T}}}m_{A,k}^{t}\end{bmatrix},$
$\Delta\tilde{\theta}_{k}^{t}\!=\!\begin{bmatrix}{\Delta\theta_{B,k}^{t\textsf{T}}}\Delta\theta_{A,k}^{t}\end{bmatrix},\tilde{\theta}_{k}^{t}\!=\!\begin{bmatrix}{\theta_{B,k}^{t\textsf{T}}}\theta_{A,k}^{t}\end{bmatrix},$ (6)
where []delimited-[][$\cdot$][ ⋅ ] is concatenation operation. With these definitions, the overall error memory update can be succinctly written as
$\tilde{m}_{k}^{t+1}=\tilde{m}_{k}^{t}+\Delta\tilde{\theta}_{k}^{t}-\tilde{\theta}_{k}^{t}.$ (7)
This process effectively corrects the errors introduced by sparsification, ensuring that the model updates are more accurate over time. The error feedback mechanism enhances the convergence of the FL process despite the sparsification. The advantage of our proposed sparsification method will be evaluated in Subsection V-A and summarized in Algorithm 1.
We state a lemma that shows the errors maintained in Algorithm 1 using the following assumption[9].
Assumption 1.
Let $\mathcal{S}:\mathbb{R}^{q}\rightarrow\mathbb{R}^{q}$ be a sparsification operator and $u$ be the number of non-zero elements satisfying $0<u<q.$ Then, for all $x\in\mathbb{R}^{q}$, the operator $\mathcal{S}$ satisfies $\mathbb{E}\left[\|\mathcal{S}(x)-x\|_{2}^{2}\right]\leq\left(1-\frac{u}{q}\right)\|x\|_{2}^{2}.$
Lemma 1.
Under the orthogonal properties, sparsification rate $O_{k}$ and Assumption 1, the expected squared norm of the sparsification error $\tilde{m}_{k}^{t+1}$ at iteration $t$ is upper-bounded as
$\mathbb{E}\left[\left\|\tilde{m}_{k}^{t+1}\right\|_{F}^{2}\right]\leq\frac{4(1-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}.$ (8)
Proof:
See Appendix A. ∎
Corollary 1.
Under the same assumptions and conditions as in Lemma 1, the sparsification error $m_{k}^{t+1}\in\mathbb{R}^{d\times\ell}$, i.e. $m_{k}^{t+1}=m_{B,k}^{t}m_{A,k}^{t}$ at iteration $t$ is upper-bounded as
$\mathbb{E}\left[\left\|m_{k}^{t+1}\right\|_{F}\right]\leq\frac{2(1-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}.$ (9)
Proof:
Applying the Arithmetic and Geometric Mean inequality (AM-GM inequality), it follows that
$2\|m_{B,k}^{t+1}m_{A,k}^{t+1}\|_{F}\leq\|m_{B,k}^{t+1}\|_{F}^{2}+\|m_{A,k}^{t+1}\|_{F}^{2}=\|\tilde{m}_{k}^{t+1}\|_{F}^{2}.$ (10)
Invoking Lemma 1, we complete the proof. ∎
IV Two-stage Federated Fine-Tuning
IV-A Convergence Analysis and Problem Formulation
In LoRA-based FL, unlike conventional FL approaches, the use of LoRA modules introduces two additional factors that must be considered in the learning process. First, as discussed in Subsection II-A, the performance of LoRA is predominantly influenced by the rank $r$. Therefore, an analysis of the impact of the selected rank for fine-tuning should be included. To quantify the original fully-finetuning model and LoRA model with rank $r$, we introduce the following assumption including an inequality.
Assumption 2.
Let $\theta_{o}$ denote the LoRA model that achieves the same performance as the fully fine-tuned model, and let $\theta_{r}$ be the LoRA model of rank $r$ trained using the loss function (III-A). Denote $\sigma(\theta_{r})$ the largest singular value of $\theta_{r}$, i.e. $\sigma(\theta_{r})=\max_{i}\|\theta_{B}[:,i]\|_{2}\|\theta_{A}[i,:]\|_{2}$, for $i=1,\dots,r$. Then, there exists a constant $H>0$ such that the LoRA error satisfies:
$\mathbb{E}[\|\theta_{o}-\theta_{r}\|_{F}^{2}]\leq H(r_{\max}-r)\sigma^{2}(\theta_{r}),$ (11)
where $r_{\max}$ is the rank of $\theta_{o}$.
Remark 1.
Assumption 2 is inspired by the SVD. Since $\theta_{r}$ is trained to maximize performance for a given rank $r$, it is expected that the dominant singular components of $\theta_{o}$ are well approximated by $\theta_{r}$. In other words, the first $r$ singular values of $\theta_{r}$ are assumed to closely match those of $\theta_{o}$. Therefore, $\|\theta_{o}-\theta_{r}\|_{F}^{2}$ can be upper bounded by $(r_{\max}-r)$ copies of the squared $r$-th largest singular value. This behavior indicates that the error between the optimal model $\theta_{o}$ and the rank $r$ LoRA model $\theta_{r}$ diminishes with increasing $r$.
Second, as mentioned in Subsection II-B, clients transmit the matrices $\theta_{B}$ and $\theta_{A}$ instead of the full matrix product $\theta_{B}\theta_{A}$. While this approach reduces computational overhead, it introduces an unexpected additional term when aggregating the updates at the server. The additional term introduced in the aggregation at the sever is as follows.
$\mathbb{E}_{k}[\theta_{B,k}\theta_{A,k}]-\theta_{B}\theta_{A}$
$=\mathbb{E}_{k}[(\theta_{B,k}-\theta_{B})](\theta_{A,k}-\theta_{A})]=\textit{Cov}(\theta_{B},\theta_{A}).$ (12)
As observed in (IV-A), $\textit{Cov}(\theta_{B},\theta_{A})$ reflects the discrepancy between the aggregated model and the individual client models. In FL, each client begins parameter refinement from a common initial model. Consequently, the discrepancy increases when the data distributions across clients are more heterogeneous [15], since the locally updated models tend to diverge in different directions, leading to an increase in $\textit{Cov}(\theta_{B},\theta_{A})$. Furthermore, the model size plays a critical role in this phenomenon. Larger models possess a higher number of parameters, which naturally leads to a greater overall parameter magnitude measured by the Frobenius norm $\|\theta_{r}\|_{F}^{2}$.
To formally account for this behavior, we introduce the following assumption:
Assumption 3.
Let $\theta_{r}$ denote the LoRA model with rank $r$. The expected Frobenius norm of the covariance of $\theta_{B}$ and $\theta_{A}$ satisfies:
$\mathbb{E}[\|\textit{Cov}(\theta_{B},\theta_{A})\|_{F}^{2}]\leq\phi\|\theta_{r}\|_{F}^{2},$ (13)
where $\phi$ is a constant that quantifies the degree of data heterogeneity.
Remark 2.
Assumption 3 indicates that the covariance between $\theta_{B}$ and $\theta_{A}$ is significantly affected by both the diversity of client data and the size of the model. Empirical evidence supporting this relationship is provided in Subsection V-C.
Next, for theoretical analysis, we make the following assumptions typically made in analyzing FL family [16].
Assumption 4.
The local objective function $F_{k}(\theta)$ is $S$-smooth, i.e. $\|\nabla F_{k}(\theta)-\nabla F_{k}(\theta^{\prime})\|_{F}\leq S\|\theta-\theta^{\prime}\|_{F},$ for all $\theta$. It also satisfies $F_{k}(\theta^{\prime})\leq F_{k}(\theta)\!+\!\nabla F_{k}(\theta)^{\textsf{T}}(\theta^{\prime}\!-\!\theta)\!+\!\frac{S}{2}\|\theta^{\prime}\!-\!\theta\|_{F}^{2},$ for all $\theta$. The global objective function $F(\theta)$, being the average of the local objectives, is then also $S$-smooth and the global objective function $F(\theta)$ is lower bounded such as $F(\theta)\geq F(\theta^{*})$ for all $\theta$.
Assumption 5.
The stochastic gradient at each client is unbiased such as $\mathbb{E}[\nabla F_{k}^{\scriptstyle{t,e}}(\theta)]\!=\!\nabla F_{k}(\theta)$ for all $k,t,e$.
Assumption 6.
The expected squared norm of the stochastic gradient and the largest singular value of LoRA model at each client are upper bounded as $\mathbb{E}[\|\nabla F_{k}^{\scriptstyle(t,e)}(\theta)\|_{F}^{2}]\leq G_{k}^{2}$ and $\sigma(\theta_{r})\leq W$, respectively. Hence, the Frobenius norm of the LoRA weight is bounded as $\mathbb{E}[\|\theta_{r}\|_{F}^{2}]\leq rW^{2}$.
Theorem 1.
Based on Assumptions 2, 3, 4, 5, 6 and Lemmas 1 and 3, the optimality gap of LoRA-based FL after $T$ global iterations is upper-bounded as
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\|\nabla F_{r}(\theta_{r}^{t})\right\|_{F}^{2}\right]\leq\frac{2}{\eta T}\left(\mathbb{E}\left[F_{r}(\theta_{r}^{0})\!-\!F^{*}\right]\right)$
$\!+\!\frac{1}{T}\sum_{t=0}^{T-1}\bigg{(}2S^{2}H(r_{\max}\!-\!r)W^{2}\!+\!4S^{2}\phi rW^{2}\!+\!\eta SE^{2}G^{2}$
$+\frac{8(N\!-\!K^{t})}{K^{t}(N\!-\!1)}\eta^{2}S^{2}E^{2}G^{2}\!+\!\frac{4N(1-O^{t})^{2}}{K^{t}{(O^{t})}^{4}}rS^{2}W^{4}$
$+\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}G^{2}\bigg{)},$ (14)
where $G^{2}=\max_{k}G_{k}^{2}$; $E$ and $\eta$ represent the number of local iterations and learning rate, respectively.
Proof:
See Appendix B. ∎
Remark 3.
Theorem 1 indicates that the convergence behavior of the LoRA-based FL is governed by several interconnected parameters: the sparsification ratios $O^{t}$ and the LoRA rank $r$. These factors are intrinsically linked, collectively influencing the trade-off between communication efficiency and learning performance. For instance, a higher sparsification ratio $O^{t}$ reduces communication overhead but may degrade convergence, while increasing the rank $r$ enhances model expressiveness at the cost of additional communication. By carefully balancing these factors, we can enhance both communication efficiency and learning effectiveness within the constraints of the LoRA-based FL environment.
Building upon the convergence analysis , our goal is to jointly optimize the key parameters to minimize the convergence upper bound. Specifically, we aim to determine the optimal LoRA rank $r$, sparsification ratios $O^{t}$, and bandwidth allocations $b_{k}$ to enhance both communication efficiency and learning performance under the constraints of limited bandwidth and training time. Our optimization problem is formulated as follows.
$\mathcal{P}_{1}$ $\min_{r,O^{t},b_{k}^{t}}\gamma^{T},$ (15a)
s.t. $\ \ r\leq r_{\max},r\in\mathbb{Z}^{+}$ (15b)
$\ \ \frac{1}{T}\sum\nolimits_{t=0}^{T-1}D^{t}\leq D_{th},$ (15c)
$\ \ O_{min}<O_{k}^{t}\leq 1,$ (15d)
$\ \ 0\leq b_{k}^{t}\leq 1,\quad\sum\nolimits_{k=1}^{N}b_{k}^{t}=1,$ (15e)
where $\gamma^{T}\!=\!\frac{1}{T}\!\sum_{t\!=\!0}^{T\!-\!1}2S^{2}H(r\!_{\max}\!-\!r)\!W^{2}\!+\!4S^{2}\phi rW^{2}\!+\!\eta SE^{2}G^{2}+\frac{8(N\!-\!K^{t})}{K^{t}(N\!-\!1)}\eta^{2}S^{2}E^{2}G^{2}\!+\!\frac{4N(1\!-\!O^{t})^{2}}{K^{t}{O^{t}}^{4}}rS^{2}W^{4}+\frac{E(E\!-\!1)(2E\!-\!1)S^{2}\eta^{2}}{6}G^{2}$, which is the upper-bound on the optimality gap in Theorem 1; (15b) is the rank feasibility constraint; (15c) is the average transmission delay constraint; (15d) is the sparsification ratio constraint with minimum sparsification ratio $O_{min}$; (15e) is the constraints for the bandwidth allocation strategy.
Note that Problem $\mathcal{P}_{1}$ presents two key challenges. First, it must be addressed prior to training, as the model structure, i.e. LoRA rank $r$, needs to be determined in advance. However, this approach would require multiple training iterations and full knowledge of the channel state information (CSI) for all clients at each iteration, which is impractical in the real scenario. Second, the presence of long-term constraints, highly coupled variables, and mixed-integer non-linear programming (MINLP) problem further makes the problem NP-hard. To address these challenges, we divide the problem into two stages. In the offline stage, we optimize the LoRA rank $r$ before training begins by analyzing the trade-off between model capacity and communication overhead. In the online stage, we adaptively manage the exact parameters except $r$ during training, based on real-time network conditions and resource availability.
IV-B Offline Stage
As discussed in the previous subsection, the offline stage is conducted prior to training with the specific goal of optimizing the appropriate LoRA rank $r$. Since the actual CSI cannot be obtained during the offline stage, the channel is approximated by assuming the average channel as the actual channel. The initial sparsification ratio are set to $O^{0}$. The offline stage thus considers the following approximated problem:
$\mathcal{P}_{2}$ $\min_{r,O^{0},b_{k}^{0}}\gamma^{0},$ (16a)
s.t. $\ \ r\leq r_{\max},~{}r\in\mathbb{Z}^{+}$ (16b)
$\ \ D^{0}\leq D_{th},$ (16c)
$\ \ O_{min}<O^{0}\leq 1,$ (16d)
$\ \ 0\leq b_{k}^{0}\leq 1,\quad\sum\nolimits_{k=1}^{N}b_{k}^{0}=1,$ (16e)
where $\gamma^{0}=2S^{2}H(r_{\max}\!-\!r)W^{2}\!+\!4S^{2}\phi rW^{2}\!+\!\eta SE^{2}G^{2}\!+\!\frac{8(N\!-\!K^{0})}{K^{0}(N\!-\!1)}\eta^{2}S^{2}E^{2}G^{2}\!+\!\frac{4N(1-O^{0})^{2}}{K^{0}{O^{0}}^{4}}rS^{2}W^{4}\!+\!\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}G^{2}$, which is the upper-bound on the initial optimality gap in Theorem 1 when $T=0$.
Note that the Problem $\mathcal{P}_{2}$ is still an NP-hard MINLP problem due to the discrete rank value and non-convexity caused by the coupled variables. To address these challenges, we first fix the integer variable $r$ to a specific value and solve the resulting continuous optimization problem $O^{0}$. By iterating over all feasible values of $r$ within the constraint, we can identify the optimal integer $r$. The overall process is in the Algorithm 2. To efficiently solve the continuous optimization problem for a fixed $r$, we present the following lemma, which provides the optimal values of $O^{0}$ given an integer $r$.
Algorithm 2 Offline Stage Optimization Algorithm
1:Initialize $\gamma^{0}_{\text{opt}}\leftarrow\infty$
2:for $r=1$ to $r_{\text{max}}$ do
3:    Fix the LoRA rank $r$
4:    Solve for $O^{0}$ using Lemma 2
5:    Compute the objective function $\gamma^{0}$ with $r$ and $O^{0}$
6:    if $\gamma^{0}<\gamma^{0}_{\text{opt}}$ then
7:       Update $\bar{r}\leftarrow r$
8:    end if
9:end for
10:return $\bar{r}$
Lemma 2.
In Problem $\mathcal{P}_{2}$, given an integer value $r$, $O^{0}$ that minimize the objective function can be obtained as
$O^{0}=\max\left(O_{min},\min\left(1,\frac{N}{AK^{0}r}\right)\right),$ (17)
where $A=\sum_{k=1}^{N}\frac{v(d+\ell)}{B\log_{2}(1+|h_{k}|^{2}/\sigma^{2})}$.
Proof:
[…]
$D_{th}$ $=\mathbb{E}\left[\sum_{k\in\mathcal{K}^{0}}\frac{vO^{0}r(d+\ell)}{B\log_{2}(1+\frac{|h_{k}|^{2}}{\sigma^{2}})}\right]$ (18)
$=\frac{K^{0}}{N}\sum_{k=1}^{N}\frac{vO^{0}r(d+\ell)}{B\log_{2}(1+\frac{|h_{k}|^{2}}{\sigma^{2}})}.$ (19)
Consequently, using the property $O_{min}<O^{0}\leq 1$, we can obtain
$O^{0}=\max\left(O_{min},\min\left(1,\frac{N}{AK^{0}r}\right)\right),$ (20)
where $A=\sum_{k=1}^{N}\frac{v(d+\ell)}{B\log_{2}(1+|h_{k}|^{2}/\sigma^{2})}$. ∎
Based on Lemma 2 and Algorithm 2, we can obtain the optimal LoRA rank to be used in training. Subsequently, clients perform fine-tuning based on the specified LoRA rank $r$. At each iteration, an online-stage optimization is conducted to determine the optimal values of $O^{t}$ and $b_{k}^{t}$.
IV-C Online Stage
In this subsection, we address the optimization of the remaining parameters for the specified LoRA rank. Our objective is to dynamically adjust the sparsification ratio $O^{t}$, and the bandwidth allocation ${b_{k}^{t}}$ for every iteration $t$. These adjustments are made in response to real-time network conditions and the availability of resources. Since these parameters are interdependent and subject to long-term latency constraints, Lyapunov optimization is employed to perform online optimization at each iteration. To handle the time-average latency constraint specified in (15c), we introduce a virtual queue $Q^{t}$ that evolves according to the following update rule
$Q^{t+1}=\max(Q^{t}+D^{t}-D_{th},0),$ (21)
where $Q^{0}\!=\!0$. The virtual queue $Q^{t}$ effectively tracks the accumulation of latency violations over time. Ensuring the stability of this queue is equivalent to satisfying the long-term latency constraint. To measure the congestion of queue $Q^{t}$, the Lyapunov function is defined as
$L(Q^{t})=\frac{1}{2}(D^{t})^{2}.$ (22)
To bound the increase of the virtual queues, we need to constrain the expected increase of the Lyapunov function. Thus the Lyapunov drift at iteration $t$ is
$\Delta(Q^{t})=\mathbb{E}[L(Q^{t+1})-L(Q^{t})|Q^{t}],$ (23)
where the expectation is taken over the random system state, i.e. channel state. To balance queue stability with the minimization of the convergence upper bound $\gamma^{t}$, we introduce the drift-plus-penalty function
$\Delta(Q^{t})+V\mathbb{E}\left[\gamma^{t}|Q^{t}\right],$ (24)
where $V>0$ is a control parameter that weights the importance of the convergence performance relative to queue stability. A larger $V$ places more emphasis on optimizing the convergence bound. To make the optimization tractable, we derive an upper bound for the Lyapunov drift. Expanding the drift expression, we have:
$\Delta(Q^{t})$
$=\mathbb{E}\!\left[\!\frac{1}{2}\!\left(\max\left\{Q^{t}\!+\!D^{t}\!-\!D_{th},0\right\}\right)^{2}\!-\!\frac{1}{2}(Q^{t})^{2}\!\Big{|}Q^{t}\!\right]$ (25)
$\leq\mathbb{E}\left[\frac{1}{2}\left(Q^{t}\!+\!D^{t}\!-\!D_{th}\right)^{2}\!-\!\frac{1}{2}(Q^{t})^{2}\Big{|}Q^{t}\right]$ (26)
$=Q^{t}\mathbb{E}\left[D^{t}\!-\!D_{th}|Q^{t}\right]\!+\!\frac{1}{2}\mathbb{E}\left[\left(D^{t}\!-\!D_{th}\right)^{2}|Q^{t}\right]$ (27)
$\leq B+Q^{t}\left(\mathbb{E}\left[D^{t}|Q^{t}\right]-D_{th}\right),$ (28)
where $B=\dfrac{1}{2}(\bar{D}-D_{th})^{2}$, and $\bar{D}$ represents the upper bound of all possible $D^{t}$. Note that $\bar{D}$ can be empirically determined, for example, by assuming that all clients receive the minimum possible bandwidth and experience the worst-case channel conditions. The inequality in (26) holds because $(\max{x,0})^{2}\leq x^{2},;\forall x\in\mathbb{R}$, while the inequality in (28) follows from the fact that $D^{t}\leq\bar{D}$.
By substituting the upper bound Lyapunov drift into the drift-plus-penalty function, we obtain an expression that we can optimize at each time slot. Note that since the current queue $Q^{t}$ is already observed at the beginning of time slot $t$, the expectation in the drift is over $D^{t}$ which affects the transition $Q^{t}$ to $Q^{t+1}$. Also, under the assumption of perfect knowledge of CSI at the server, the channel randomness at time $t$ is fully observed at the beginning of the time slot. Consequently, both $D^{t}$ and $\gamma^{t}$ become deterministic functions of the observed CSI at time $t$. Discarding constant terms that do not affect the optimization, we reformulate the per-slot problem based on the drift-plus-penalty function as follows:
$\mathcal{P}_{3}:$ $\min_{O^{t},b_{k}^{t}}Q^{t}D^{t}+V\gamma^{t}$ (29a)
s.t. $\ \ O_{\min}\leq O^{t}\leq 1,$ (29b)
$\ \ 0\leq b_{k}^{t}\leq 1,\quad\sum\nolimits_{k=1}^{N}b_{k}^{t}=1,$ (29c)
where $\gamma^{t}=2S^{2}H(r_{\max}\!-\!r)W^{2}\!+\!4S^{2}\phi rW^{2}\!+\!\eta SE^{2}G^{2}+\frac{8(N\!-\!K^{t})}{K^{t}(N\!-\!1)}\eta^{2}S^{2}E^{2}G^{2}\!+\!\frac{4N(1-O^{t})^{2}}{K^{t}{O^{t}}^{4}}rS^{2}W^{4}\!+\!\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}G^{2}$, which is the upper-bound on the optimaility gap at time $t$ in Theorem 1.
To address Problem $\mathcal{P}_{3}$, we begin by demonstrating that it is convex with respect to $O^{t}$. Specifically, the second derivative of the objective function is given by
$\frac{8NrS^{2}W^{2}}{K^{t}{O^{t}}^{6}}(3{O^{t}}^{2}-12O^{t}+10),$ (30)
which is always positive due to the constraint $O_{\text{min}}<O^{t}\leq 1$. This confirms the convexity of the problem, allowing us to employ efficient convex optimization algorithms or well-established software tools such as CVX to find the global optimum.
For determining the bandwidth allocation strategy $b_{k}^{t}$, we leverage the property that only selected clients, i.e. those with $a_{k}^{t}=1$, receive nonzero bandwidth, i.e. $b_{k}^{t}>0$, which allows us to derive
$b_{k}^{t}\!=\!\frac{vO^{t}r(d\!+\!\ell)}{\!BD^{t}\!\log_{2}\!\left(\!1\!+\!\frac{|h_{k}|^{2}}{\sigma^{2}}\!\right)}\!=\!\frac{1}{\!A^{\prime}\!\log_{2}\!\left(\!1\!+\!\frac{|h_{k}|^{2}}{\sigma^{2}}\!\right)},$ (31)
where $A^{\prime}=\sum_{k\in\mathcal{K}^{t}}\frac{1}{\log_{2}\left(1+\frac{|h_{k}|^{2}}{\sigma^{2}}\right)}$. This expression indicates that the bandwidth allocation for each client is inversely proportional to the logarithm of their channel gain-to-noise ratio, promoting fairness among clients with varying channel conditions.
At each iteration, the server computes the optimal values of $b_{k}^{t}$ and $O^{t}$ and communicates these parameters to the selected clients. The clients then utilize the allocated bandwidth and parameters for data transmission. The overall procedure of the proposed WFLoRA algorithm is summarized in Algorithm 3.
Algorithm 3 Two Stage Federated Algorithm
1:Offline Stage
2:Obtain optimal LoRA rank $\bar{r}$ using Algorithm 2
3:Server broadcasts pre-trained model with LoRA rank $\bar{r}$ to all clients
4:Initialize virtual queue $Q^{0}\leftarrow 0$
5:Online Stage
6:for $t=0$ to $T-1$ do
7:    At the server:
8:     Observe current network conditions and $Q^{t}$
9:     Solve Problem $\mathcal{P}_{3}$ to obtain optimal $O^{t}$ and $b_{k}^{t}$
10:     Broadcast $O^{t}$, $b_{k}^{t}$, $\theta_{B}^{t}$, $\theta_{A}^{t}$ to selected clients
11:    At each selected client $k\in\mathcal{K}^{t}$:
12:     Initialize local LoRA matrices $\theta_{B,k}^{t}\leftarrow\theta_{B}^{t}$, $\theta_{A,k}^{t}\leftarrow\theta_{A}^{t}$
13:     Perform local training using loss function (III-A)
14:     Apply SOFT and transmit $\theta_{B,k}^{t}$ and $\theta_{A,k}^{t}$ to the server
15:    At the server:
16:     Aggregate received updates:
17:     $\theta_{B}^{t+1}\leftarrow\dfrac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{B,k}^{t}$
18:     $\theta_{A}^{t+1}\leftarrow\dfrac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{A,k}^{t}$
19:     Update global model: $\theta_{r}^{t+1}\leftarrow\theta_{P}+\dfrac{\alpha}{r}\theta_{B}^{t+1}\theta_{A}^{t+1}$
20:     Update virtual queue: $Q^{t+1}\leftarrow\max(Q^{t}+D^{t}-D_{th},0)$
21:end for
V Numerical Results
In this section, we present simulation results to demonstrate the effectiveness of the proposed SOFT and TSFA framework. All experiments were conducted using Python 3.8 on an Ubuntu server equipped with NVIDIA GeForce RTX 3090 GPUs. The total number of clients is $100$ in all experiments.
We evaluated the performance of our method on the CIFAR-100 datasets. The CIFAR-100 dataset consists of $60,000$ with $32\times 32$ color images in $100$ classes, with $600$ images per class. There are $50,000$ in training images and $10,000$ in test images. The classes are grouped into $20$ superclasses, each containing five classes, covering a wide range of objects such as animals, vehicles, and everyday items. In our experiments, we utilized transformer-based pre-trained models for fine-tuning. Specifically, we employed the ViT-Base model, which stands for the Vision Transformer Base model. ViT is a transformer-based architecture specifically designed for image recognition tasks.
Regarding the baseline schemes in the simulation, we consider the following frameworks:
  • Top-q LoRA (TLoRA)[17, 9]: Each client selects and transmits only the elements of the LoRA matrices with the largest magnitudes. Specifically, after computing the low-rank adaptation matrices, clients identify the top-$q$ elements with the highest absolute values and sparsify the rest.
  • Random LoRA (RLoRA)[9]: Each client selects and transmits only the elements of the LoRA matrices randomly without considering their magnitudes or positions within the matrices.
  • Structured LoRA (SLoRA)[11]: Each client selects and transmits only the elements of the LoRA matrices in a structured manner. Specifically, each client preserves the elements corresponding to lower rank indices and applies sparsification to those associated with higher rank indices. Here, rank indice refers to the ordering of LoRA components as defined by the model’s architecture.
  • Ideal LoRA (ILoRA)[3]: The Ideal LoRA framework represents an idealized environment where there are no communication constraints between clients and the server. In this scenario, clients transmit full LoRA matrices without any sparsification. This serves as an upper bound for performance.
V-A Effects of SOFT
(a) LoRA rank $r=4$
(b) LoRA rank $r=8$
(c) OSFA and TSFA
Figure 3: Test accuracy of CIFAR-100 classification. (a) and (b) show the performance using LoRA with ranks $r=4,r=8$, respectively, under various sparsification methods, while panel (c) compares the performance of OSFA and TSFA.
In this subsection, we evaluate the effectiveness of the SOFT method by comparing it with the baseline methods. Figs. 3(a) and 3(b) illustrate the test accuracy on the CIFAR-100 dataset using various sparsification methods with LoRA rank $r=4$ and $r=8$, respectively. We set the sparsification ratio as $0.5$ for all baselines and the proposed schemes. Notably, SOFT consistently maintains higher accuracy compared to the other schemes. This superior performance can be attributed to SOFT’s ability to dynamically sparsify the LoRA modules by considering the importance of individual parameters. In LoRA module-based fine-tuning, two matrices are multiplied to form a single layer. TLoRA exhibits lower performance in this context. This is because simply selecting the top-$k$ elements does not effectively capture the crucial interactions between the two matrices involved in LoRA. As a result, the essential information required for accurate model updates may be lost, leading to suboptimal performance. Similarly, SLoRA, which transmits only the elements associated with the lower rank indices throughout training, underperforms SOFT. Since SLoRA sets a fixed sparsity pattern without adapting to the varying importance of parameters during training, it cannot capture the dynamic changes in the model’s weight distribution. This rigidity results in less effective learning and lower accuracy. Furthermore, RLoRA is shown to be unsuitable for LoRA-based fine-tuning. Randomly selecting parameters to transmit fails to account for their significance in the model, which can hinder the learning process and prevent convergence. This approach leads to poor model performance.
V-B Effects of Offline and Online Stage
In Section IV, we present an optimization framework that employs a two stage process: an offline stage for pre-training configuration and an online stage for dynamic parameter adjustment during training. In this subsection, we evaluate the performance of TSFA by comparing it with a one-stage variant (OSFA), which forgoes pre-optimization of the LoRA rank $r$ and instead relies solely on online adjustments with arbitrarily chosen $r$ values. The experiments were conducted using $N=100$, $K^{t}=10$, and $V=0.0001$ under a Rayleigh block fading channel characterized by a zero mean and unit variance, with constant channel gains during each user uplink transmission. The proposed SOFT method was utilized for sparsification.
Fig. 3(c) clearly demonstrates that pre-selecting an appropriate LoRA rank $r$ is essential for optimal performance. When OSFA with $r=1$, $r=2$ or $r=4$, the model suffers from insufficient capacity, despite benefiting from lower sparsification ratios compared to TSFA. In these cases, the reduced rank is inadequate to capture the model’s complexity, leading to under-parameterization and poor learning of the underlying data distribution. Conversely, when OSFA employs higher $r$ values such as $r=16$, $r=32$, or $r=64$, the increased model capacity requires a corresponding increase in the sparsification ratio to satisfy communication constraints. This, in turn, results in excessive sparsification, which limits the transmission of critical parameter updates and adversely affects learning. These findings underscore the importance of the offline stage in TSFA. By determining an appropriate LoRA rank $r$ prior to training, the two stage approach effectively balances model capacity and communication efficiency, thereby avoiding the pitfalls of both under-parameterization and over-sparsification.
V-C Effects of Covariance
Figure 4: Frobenius norm of covariance under different number of shards.
Figure 5: Test accuracy of CIFAR-100 under different number of shards.
In Section II, we introduced the covariance term that naturally arises due to the separate transmission and aggregation of LoRA matrices. In this subsection, we provide empirical evidence illustrating the strong correlation between this covariance term and the degree of data heterogeneity in a non-IID setting. To simulate a heterogeneous environment, we employ a shard-based non-IID partitioning strategy. We begin by sorting the dataset by class labels, ensuring that each subgroup is composed of samples belonging to the same class. These sorted subsets are then split into a predetermined number of shards, with each shard predominantly containing data from a single class. By randomly distributing these shards to clients, each client receives data biased toward certain classes, thereby introducing data heterogeneity.
Fig. 4 shows how the covariance term evolves during training for different shard sizes, and Fig. 5 shows the corresponding test accuracy. We set $r=32$ and assume no communication constraint in order to sorely consider non-IID effects. As the shard size increases, the sampling process becomes increasingly similar to an IID scenario. Larger shards yield more balanced client data distributions, which naturally reduce the discrepancy among locally updated LoRA parameters. This results in a lower covariance magnitude, as evidenced by the decreasing trend in Fig. 4. The results in Fig. 5 confirm this relationship, scenarios with lower covariance consistently achieve higher test accuracy, highlighting the pivotal role that covariance plays in model convergence and generalization under non-IID conditions. Developing methods to reduce or compensate for this effect whether through data-driven approaches, adaptive averaging mechanisms, or improved LoRA configurations, remains an intriguing avenue for future research.
VI Conclusion
In this paper, we proposed a wireless federated LoRA fine-tuning framework that bridges the gap between state-of-the-art LLMs and the practical constraints of FL environments. Our analysis underscores the impact of LoRA rank and covariance on learning dynamics. We introduced SOFT for efficient parameter selection and TSFA for dynamically optimizing sparsification and bandwidth. By leveraging Lyapunov optimization, TSFA ensures long-term latency stability and resource efficiency. Experiments show that the proposed framework, enhanced with SOFT and TSFA, achieves performance on par with centralized methods while significantly reducing communication overhead, providing a scalable solution for large-scale FL over wireless networks.
Appendix A Proof of Lemma 1
Before proceeding with the main proof, we first establish essential inequality. Using Arithmetic and Geometric Mean inequality (AM-GM inequality), i.e. $\gamma a+\frac{1}{\gamma}b\geq 2\sqrt{ab}$ for $\gamma>0$, $\|x+y\|_{F}^{2}$ can be upper-bounded as:
$\|x+y\|_{F}^{2}$ $=\|x\|_{F}^{2}+\|y\|_{F}^{2}+2x^{\textsf{T}}y$ (32)
$\leq\|x\|_{F}^{2}+\|y\|_{F}^{2}+\gamma\|x\|_{F}^{2}+\frac{1}{\gamma}\|y\|_{F}^{2}$ (33)
$=(1+\gamma)\|x\|_{F}^{2}+\left(1+\frac{1}{\gamma}\right)\|y\|_{F}^{2}.$ (34)
Now, using the Assumption 1 and (7),
$\mathbb{E}\left[\left\|\tilde{m}_{k}^{t+1}\right\|_{F}^{2}\right]$ $\leq(1-O_{k})\mathbb{E}\left[\|\tilde{m}_{k}^{t}+\Delta\tilde{\theta}_{k}^{t}\|_{F}^{2}\right]$ (35)
$\leq(1-O_{k})(1+\gamma)\mathbb{E}[\|\tilde{m}_{k}^{t}\|_{F}^{2}]$
$\quad+(1-O_{k})\left(1+\frac{1}{\gamma}\right)\|\Delta\tilde{\theta}_{k}^{t}\|_{F}^{2},$ (36)
where the second inequality just follows (34). Finally, using (36) recursively and define $(1-O_{k})\left(1\!+\!\frac{1}{\gamma}\right)=\Gamma$,
$\mathbb{E}[\|m_{k}^{t+1}\|_{F}^{2}]$
$\leq\Gamma\sum\nolimits_{i=0}^{t}\left[(1\!-\!O_{k})(1\!+\!\gamma)\right]^{t\!-\!i}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ (37)
$\leq\Gamma\sum\nolimits_{j=0}^{t}\left[(1-O_{k})(1+\gamma)\right]^{j}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ (38)
$\leq\Gamma\sum\nolimits_{j=0}^{\infty}\left[(1-O_{k})(1+\gamma)\right]^{j}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ (39)
$=\frac{2(1-O_{k})(2-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ (40)
$\leq\frac{4(1-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2},$ (41)
where (38) is by $\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}\leq\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ and (40) is satisfied by $\gamma=\frac{O_{k}}{2(1-O_{k})}$, which ensures $(1+\gamma)(1-O_{k})<1$ for $10<O_{k}<$ 1. Moreover, the last inequality is due to $2-O_{k}\leq 2$.
Appendix B Proof of Theorem 1
Let $\theta_{o}$ be an original model. Then the update of the original and LoRA model at the $t+1$-th global iteration can be described as:
$\theta_{o}^{t+1}$ $=\sum\nolimits_{k\in\mathcal{N}}p_{k}\theta_{o,k}^{t+1}$
$=\theta_{o}^{t}\!-\!\sum\nolimits_{k\in\mathcal{N}}p_{k}\eta\sum_{e=0}^{E-1}\nabla F_{o,k}(\theta_{o,k}^{t,e}),$ (42)
$\theta_{r}^{t+1}$ $=\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{r,k}^{t+1}\!+\!p_{k}m_{k}^{t}+\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1})$
$=\theta_{r}^{t}\!-\!\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\eta\sum_{e=0}^{E-1}\nabla F_{r,k}(\theta_{r,k}^{t,e})\!+\!p_{k}m_{k}^{t}$
$\quad\quad+\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1}),$ (43)
where $m_{k}^{t}=m_{B}^{t}m_{A}^{t}$.
Using $S$-smoothness,
$\mathbb{E}[F_{o}(\theta_{o}^{t+1})\!-\!F_{o}(\theta_{o}^{t})]$ $\leq\underbrace{\mathbb{E}[\nabla F_{o}(\theta_{o}^{t})^{\textsf{T}}(\theta_{o}^{t+1}\!-\!\theta_{o}^{t})]}_{(a)}$
$+\!\underbrace{\frac{S}{2}\mathbb{E}[\|\theta_{o}^{t+1}\!-\!\theta_{o}^{t}\|_{F}^{2}]}_{(b)}.$ (44)
For $(a)$, using (42), we can obtain
$(a)$ $=-\eta\sum_{e=0}^{E-1}\mathbb{E}\left[\nabla F_{o}(\theta_{o}^{t})^{\textsf{T}}\sum_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o}^{t,e})\right]$ (45)
$=-\frac{\eta}{2}\underbrace{\sum\nolimits_{e=0}^{E-1}\mathbb{E}\left[\left\|\nabla F_{o}(\theta_{o}^{t})\right\|_{F}^{2}\right]}_{(a1)}$
$-\frac{\eta}{2}\underbrace{\sum\nolimits_{e=0}^{E-1}\mathbb{E}\left[\left\|\sum\nolimits_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o,k}^{t,e})\right\|_{F}^{2}\right]}_{(a2)}$
$\!+\!\frac{\eta}{2}\underbrace{\sum_{e=0}^{E-1}\!\mathbb{E}\!\left[\!\left\|\!\nabla\!F_{o}(\theta_{o}^{t})\!-\!\sum_{k\in\mathcal{N}}p_{k}\!\nabla\!F_{o,k}(\theta_{o,k}^{t,e})\!\right\|_{F}^{2}\!\right]}_{(a3)},$ (46)
where the second equality is due to the property $-\mathbf{a}^{\textsf{T}}\mathbf{b}=\frac{1}{2}(-\|\mathbf{a}\|_{2}^{2}-\|\mathbf{b}\|_{2}^{2}+\|\mathbf{a-b}\|_{2}^{2})$. Next, we use the following lemma:
Lemma 3.
Under Assumptions 2, 3, and 6, the inequality below is satisfied:
$-\sum_{e=0}^{E-1}\mathbb{E}\left[\left\|\sum_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o,k}^{t,e})\right\|_{F}^{2}\right]$
$\leq\!-\!\mathbb{E}[\|\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}]\!+\!2S^{2}H(r\!_{\max}\!-\!r)W^{2}\!+\!4S^{2}\phi rW^{2}$
$\;\;\!+\!\frac{8(N\!-\!K^{t})}{K^{t}(N\!-\!1)}\!\eta^{2}S^{2}E^{2}G^{2}\!+\!\frac{\!4N(1\!-\!O^{t})\!^{2}}{K^{t}{O^{t}}^{4}}rS^{2}W^{4}.$ (47)
Proof:
See Appendix C. ∎
Using Lemma 3, we can derive an upper-bound on $(a2)$. Next, using the $S$-smoothness, Jensen’s inequality and the properties $\theta_{o}^{t}=\theta_{o,k}^{t,0}$ and $\sum_{k\in\mathcal{N}}p_{k}=1$, we can obtain an upper-bound on $(a3)$ as:
$(a3)\!$ $\leq S^{2}\sum_{k\in\mathcal{N}}p_{k}\sum_{e=0}^{E-1}\mathbb{E}\left[\left\|\theta_{o,k}^{t,0}-\theta_{o,k}^{t,e}\right\|_{F}^{2}\right]$ (48)
$=S^{2}\sum_{k\in\mathcal{N}}p_{k}\sum_{e=0}^{E-1}\!\mathbb{E}\!\left[\!\left\|\sum_{i=0}^{e-1}\!\eta\nabla F_{o,k}(\theta_{o,k}^{t,e})\!\right\|_{F}^{2}\!\right]$ (49)
$\leq S^{2}\eta^{2}\sum_{k\in\mathcal{N}}p_{k}\sum_{e=0}^{E-1}e^{2}G_{k}^{2}$ (50)
$=\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}\sum_{k\in\mathcal{N}}p_{k}G_{k}^{2}$ (51)
$\leq\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}G^{2},$ (52)
where the first equality comes from (42); the first inequality is due to Jensen’s inequality and Assumption 6; the last inequality is by $\!G^{2}\!=\!\max_{k}\!G_{k}^{2}\!$ and $\!\sum_{k\in\mathcal{N}}\!p_{k}\!\!=\!\!1$.
Similarly, using (43) and Jensen’s inequality, $(b)$ in (B) can be upper-bounded as:
$(b)$ $\leq\eta\mathbb{E}\left[\sum_{k\in\mathcal{N}}p_{k}\sum_{e=0}^{E-1}\left\|\nabla F_{o,k}^{t,e}(\theta_{o}^{t,e})\right\|_{F}^{2}\right]$ (53)
$\leq\eta^{2}E^{2}\sum_{k\in\mathcal{N}}p_{k}G_{k}^{2},$ (54)
$\leq\eta^{2}E^{2}G^{2},$ (55)
where the second inequality comes from Assumption 6. The last inequality is due to $G^{2}=\max_{k}{G_{k}^{2}}$ and $\sum_{k\in\mathcal{N}}p_{k}=1$.
Finally, applying the above inequalities to (B) and using $(a1)>0$ always holds, we can obtain
$\mathbb{E}[F_{o}\!(\theta_{o}^{t+1}\!)\!-\!F_{o}\!(\theta_{o}^{t}\!)]\!\leq\!-\!\frac{\eta}{2}\mathbb{E}[\|\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}]\!+\!\frac{S}{2}\!\eta^{2}\!E^{2}\!G^{2}$
$\!+\!\eta(r_{\!\max}\!-\!r)\!H\!S^{2}\!W^{2}\!\!+\!2\eta S^{2}\!\phi rW^{2}\!\!+\!\frac{\!4\eta^{3}\!(N\!-\!K^{t}\!)}{K^{t}(N\!-\!1)}\!S^{2}\!E^{2}\!G^{2}$
$\!+\!\frac{2\eta N}{\!K^{t}(O^{t})^{4}}\!(1\!-\!O^{t})^{2}rS^{2}\!W^{4}\!+\!\frac{\!\eta^{3}\!E(E\!-\!1)\!(2E\!-\!1)}{12}\!S^{2}\!G^{2}.$ (56)
Averaging the above inequality over iteration from 00 to $T-1$ and using the property that $\theta_{o}^{0}=\theta_{r}^{0}$ since all LoRA modules are initialized with 0 at the initial stage, we can finally obtain the upper bound in Theorem 1.
Appendix C Proof of Lemma 3
First, we choose only $e=0$ from the summation of positive norm values, then we obtain
$\sum_{e=0}^{E-1}$ $\mathbb{E}\left[\left\|\sum_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o,k}^{t,e})\right\|_{F}^{2}\right]$ (57)
$\!\geq\!\mathbb{E}\!\!\left[\!\left\|\!\sum_{k\in\mathcal{N}}\!p_{k}\!\nabla\!F_{o,k}\!(\theta_{o,k}^{t,0})\!\right\|_{F}^{2}\!\right]\!\!\!=\!\mathbb{E}[\|\nabla\!F\!_{o}(\theta_{o}^{t})\|_{F}^{2}],$ (58)
where the equality is by the initialization of the local models, $\theta_{o,k}^{t,0}=\theta_{o}^{t}$ for every client $k$. Next, we obtain $\nabla F_{r}(\theta_{r}^{t})$ by projecting $\nabla F_{o}(\theta_{o}^{t})$ onto the subspace spanned by the top $r$ singular vectors of $\theta_{o}^{t}$, thereby discarding the gradient components corresponding to the lower singular values. Since $\theta_{r}^{t}$ lies entirely within this subspace, evaluating the gradient of $F_{o}$ at $\theta_{r}^{t}$ is equivalent to evaluating the gradient of the restricted function $F_{r}$. Using a simple property, $\|\nabla F_{o}(\theta_{o}^{t})\|_{F}^{2}\geq\|\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}$, we have
$-\mathbb{E}[\|\nabla$ $F_{o}(\theta_{o}^{t})\|_{F}^{2}]\leq-\mathbb{E}[\|\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}]$
$+\underbrace{\mathbb{E}[\|\nabla F_{o}(\theta_{o}^{t})-\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}]}_{(c)}.$ (59)
Then, from $S$-smoothness, $(c)$ is upper bounded as
$(c)\!$ $\leq\!S^{2}\mathbb{E}[\|\theta_{o}^{t}-\theta_{r}^{t}\|_{F}^{2}]$ (60)
$\leq\!2S^{2}\underbrace{\mathbb{E}\!\left[\!\left\|\theta_{o}^{t}\!-\!\hat{\theta}_{r}^{t}\!\right\|_{F}^{2}\!\right]}_{(c1)}\!+\!2S^{2}\underbrace{\mathbb{E}\!\left[\!\left\|\hat{\theta}_{r}^{t}\!-\!\theta_{r}^{t}\right\|_{F}^{2}\!\right]}_{(c2)}$ (61)
where the second inequality is due to $\|a-b\|_{F}^{2}\leq 2\|a\|_{F}^{2}+2\|b\|_{F}^{2}$ and $\hat{\theta}_{r}^{t}=\sum_{k\in\mathcal{N}}p_{k}\theta_{B,k}^{t}\theta_{A,k}^{t}$. For $(c1)$,
$(c1)$ $\leq\mathbb{E}\left[\sum_{k\in\mathcal{N}}p_{k}\left\|\theta_{o,k}^{t+1}-\hat{\theta}_{r,k}^{t+1}\right\|_{F}^{2}\right]$ (62)
$\leq H(r_{\max}-r)W^{2},$ (63)
where the first inequality comes from Jensen’s inequality and the second inequality comes from Assumption 2 and Assumption 6. Next, plugging (43), $(c2)$ can be expressed as:
$(c2)$ $=\underbrace{\mathbb{E}\left[\left\|\sum_{k\in\mathcal{N}}p_{k}\theta_{r,k}^{t}-\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{r,k}^{t}\right\|_{F}^{2}\right]}_{(c21)}$
$+\underbrace{\mathbb{E}\left[\left\|\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}m_{k}^{t}-\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1})\right\|_{F}^{2}\right]}_{(c22)}$
$+2\bigg{\langle}\sum_{k\in\mathcal{N}}p_{k}\theta_{r,k}^{t}-\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{r,k}^{t},$
$\quad\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}m_{k}^{t}-\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1})\bigg{\rangle},$ (64)
where the last term is zero due to unbiasedness.
We follow the same steps as in Lemma 5 of [16] to derive the upper-bound $(c21)$, which utilize the probabilistic sampling properties, model unbiasedness and $\mathbb{E}[\|x-\mathbb{E}[x]\|_{F}^{2}]\leq\mathbb{E}[\|x\|_{F}^{2}]$.
$(c21)\leq\frac{4(N-K^{t})}{K^{t}(N-1)}\eta^{2}E^{2}G^{2},$ (65)
where $G^{2}=\max_{k}G_{k}^{2}$. Next, we use a simple inequality of $\|\mathbf{a}-\mathbf{b}\|_{F}^{2}\leq 2\|\mathbf{a}\|_{F}^{2}+2\|\mathbf{b}\|_{F}^{2}$, Then, $(c22)$ can be bounded as
$(c22)$ $\leq 2\mathbb{E}\left[\left\|\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}m_{k}^{t}\right\|_{F}^{2}\right]$
$+2\mathbb{E}\left[\left\|\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1})\right\|_{F}^{2}\right]$ (66)
$\leq\frac{2N}{K^{t}}\mathbb{E}\left[\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\|m_{k}^{t}\|_{F}^{2}\right]+2\phi rW^{2}$ (67)
$\leq\frac{4N(1-O^{t})^{2}}{K^{t}{(O^{t})}^{4}}rW^{4}+2\phi rW^{2}$ (68)
where the second inequality comes from Jensen’s inequality and Assumption 3; the last inequality comes from Corollary 1. Note that for the unbiasedness update, we set the sparsification ratio to be equal across the clients.