### v2

# Two-Stage Wireless Federated LoRA Fine-Tuning with Sparsified Orthogonal Updates

Abstract
Federated fine-tuning of pre-trained models with Low-Rank Adaptation (LoRA) has emerged as a promising approach for privacy-preserving on-device adaptation. However, in wireless federated settings, existing methods still face two major challenges: selecting an appropriate structural LoRA rank under communication constraints and sparsifying LoRA updates without destroying the inherent structure between the two LoRA factors. In particular, naively applying conventional sparsification to the two LoRA factors incurs severe performance loss, while repeated full matrix multiplication and singular value decomposition are computationally prohibitive for resource-constrained clients. To address these issues, we propose a wireless federated LoRA fine-tuning framework with Sparsified Orthogonal Fine-Tuning (SOFT) and a Two-Stage Federated Algorithm (TSFA). SOFT imposes near orthogonality on the LoRA factors and uses a per-rank norm-based importance metric to perform efficient SVD-free sparsification. TSFA first determines the structural LoRA rank in an offline stage and then adaptively optimizes the sparsification ratio and bandwidth allocation online via Lyapunov optimization under long-term latency constraints. We further present a convergence analysis that explicitly characterizes the effects of structural LoRA rank, sparsification error, and the covariance induced by separate aggregation of LoRA factors.
Index Terms:
Parameter-efficient fine-tuning, federated learning, sparsification, Lyapunov optimization.
I Introduction
Transformer-based large language models (LLMs) like ChatGPT [1] have demonstrated exceptional performance across a multitude of tasks, including applications in Natural Language Processing (NLP), Computer Vision (CV), Healthcare, and Biomedical research. Despite their impressive capabilities, training these models from scratch is highly resource-intensive, demanding substantial time and computational power. Consequently, it has become common practice to fine-tune pre-trained models to cater to specific domains, maximizing efficiency while leveraging existing architectures.
As a promising solution to fine-tuning, parameter-efficient fine-tuning (PEFT) methods have been developed [2, 3]. PEFT techniques focus on adjusting a smaller subset of model parameters, thereby reducing computational load and memory requirements while achieving performance comparable to that of fully fine-tuned models. Among these methods, Low-Rank Adaptation (LoRA) [3] stands out as a particularly effective approach. LoRA minimizes the number of trainable parameters by introducing trainable low-rank matrices integrated with the existing pre-trained model weights, which remain frozen during fine-tuning. This technique allows for efficient fine-tuning by updating only these additional low-rank matrices, significantly reducing computational and memory demands without substantially compromising performance. On many tasks, LoRA has been shown to achieve comparable or even superior performance compared to fully fine-tuning.
In many practical scenarios, the data necessary for fine-tuning is distributed across multiple clients, and centralizing this data raises significant privacy and security concerns. To address these challenges, federated learning (FL) has emerged as a promising paradigm for collaborative model training without the need for centralized data aggregation [4, 5]. FL enables multiple clients to jointly train a global model by sharing model updates. Each client trains the model locally using its own data and periodically communicates the model parameters or gradients to a central server, which aggregates them to update the global model. In practical wireless FL systems, however, communication efficiency becomes critical due to limited bandwidth and fluctuating channel conditions [6, 7]. To mitigate these challenges, sparsification techniques are commonly employed in FL by transmitting only the most significant model updates, typically by selecting the top-$q$ norm elements [8].
Several recent studies have attempted to combine LoRA with FL to harness the benefits of both methodologies [9, 10, 11, 12, 13, 14, 15]. For instance, the authors in [9] pioneered the integration of LoRA-based local updates with the Federated Averaging (FedAvg) algorithm [4] for model aggregation. Building upon this, [10] introduced Federated Freeze A LoRA (FFA-LoRA), a method that addresses data heterogeneity and provides privacy guarantees through differential privacy. FFA-LoRA uses only one of the two LoRA matrices during training and employs the Gaussian mechanism to analyze differential privacy. The authors in [11] investigated the impact of data heterogeneity on learning efficiency and introduced a data-driven initialization method to address this challenge. In addition, [12] proposed a personalized LoRA aggregation framework in which each client adaptively fuses the global and local LoRA parameters element-wise before local training, improving performance under heterogeneous data distributions. The authors in [13, 14] proposed a split federated LoRA framework, where the computationally intensive encoder is deployed at the edge server, while the embedding and task modules remain at the clients to account for their limited computational capabilities. Meanwhile, [15] combined LoRA with dropout-based aggregation to mitigate overfitting during training.
Despite these advances, an important problem remains unresolved. Most existing studies fix a pre-defined LoRA rank prior to training and do not offer a principled mechanism to select it under performance and resource constraints. Although some prior works [16, 17, 18] consider adaptive rank strategies, they typically operate within a pre-defined rank budget rather than determining the structural LoRA rank itself in a principled and communication-aware manner. Since performance is highly sensitive to the choice of rank r\!r\![3], this limitation can lead to either degraded model performance or inefficient resource utilization. Specifically, a rank that is too low may fail to capture essential features, whereas a higher rank increases the communication payload between clients and the server.
Moreover, most existing methods do not explicitly address time varying wireless communication conditions. Although some prior works [19, 18] attepted to reduce communication overhead through top-$q$ sparsification and uniform quantization, they applied these techniques directly to LoRA without considering its structural properties, where each update is represented by two coupled low-rank matrices. Specifically, in LoRA, model updates are factorized into two low-rank matrices. Naively applying top-$q$ selection to each matrix independently ignores the interdependence between them: small entries in one matrix may exert a disproportionately large influence when multiplied by large entries in the other. Accurately identifying the most influential updates would therefore require computing the full matrix product, followed by singular value decomposition (SVD), and transmitting only the components associated with the largest singular values. However, performing frequent SVD operations is computationally prohibitive, particularly for resource-constrained clients. As a result, both naive top-$q$ approaches and SVD-based methods become impractical in LoRA-based FL. This limitation highlights the need for novel sparsification strategies that can effectively capture the most significant information while maintaining computational efficiency.
To address these challenges, we propose a practical wireless federated LoRA fine-tuning framework tailored for resource-constrained clients. Our framework first selects a structural LoRA rank offline in a principled way and then adapts the sparsification ratio and bandwidth allocation online during training. Also, we introduce a novel LoRA sparsification strategy, sparsified orthogonal fine-tuning (SOFT), which leverages near orthogonality constraints on the LoRA matrices, without requiring full matrix multiplications or SVD operations. In this scheme, each rank vector, comprising the corresponding column vectors in the left LoRA matrix and row vectors in the right LoRA matrix, is near orthogonal. As a result, the first LoRA matrix effectively functions similarly to the left singular matrix in an SVD, while the second corresponds to the right singular matrix. To determine the importance of the $i\in\{1,...,r\}$-th rank vector, we compute the product of the norm of its column in the left matrix and the norm of its row in the right matrix, using this result as a proxy for the singular value. By enforcing these near orthogonality constraints, SOFT efficiently selects the most significant updates while maintaining low computational overhead.
On the other hand, to determine LoRA rank, we perform a convergence analysis of LoRA within the FL setting. By incorporating the effects of structural LoRA rank, sparsification, and the discrepancy induced by separate aggregation of the two LoRA factors, we derive theoretical insights into their impact on model performance and convergence behavior. Based on this analysis, we first divide the optimization parameters into two groups: parameters that must be determined before the training begins and parameters that can be adjusted dynamically during the training. Specifically, since the LoRA rank represents the model structure used in FL training which must be established prior to the training, we determine it through an offline stage. After training commences, we adaptively adjust the sparsification ratios and bandwidth allocation strategy in response to real-time network conditions and resource availability in an online stage. To manage long-term latency constraints and ensure stable convergence, we employ Lyapunov optimization.
Contributions.  The key contributions of this article are summarized as follows:
  • We develop SOFT, a LoRA sparsification method that encourages near orthogonality in the LoRA matrices via orthogonal regularization and employs a per-rank norm-based score as an SVD-free importance proxy. This enables efficient update selection without explicitly forming full product matrix or repeatedly performing SVD.
  • We provide a convergence analysis for federated LoRA fine-tuning that explicitly captures the effects of the structural LoRA rank, sparsification error, and the discrepancy induced by separate aggregation of the two LoRA factors. The analysis clarifies how structural rank selection influences approximation quality, communication efficiency, and convergence behavior.
  • We propose a two-stage federated algorithm (TSFA) framework that first determines the structural LoRA rank under approximate channel conditions before the training begins since it decides the model structure, and then adaptively adjusts the sparsification ratio and bandwidth allocation at each iteration using Lyapunov optimization.
  • We rigorously validate the effectiveness of the proposed framework through extensive simulations and experiments on benchmark datasets. The results demonstrate that our approach achieves comparable or superior performance to the existing methods while significantly reducing communication overhead, thereby substantiating its practicality and efficiency.
Organization.  The remainder of this article is organized as follows. In Section II, we describe the system model of the proposed framework, including the LoRA mechanism, FL algorithm, and communication model. Section III introduces the sparsification method. In Section IV, we present the offline and online optimization approach for parameter selection. Section V provides numerical results demonstrating the effectiveness of our method. Finally, Section VI concludes the paper. Table I provides a summary of the key notations used.
II System Model
TABLE I: Notations and Descriptions
Notation Description
$\theta_{P}$ Pre-trained model weight matrix in $\mathbb{R}^{d\times\ell}$
$p_{k}$ Data proportion for client $k$, $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$
$\mathcal{D}_{k}$ Local dataset of client $k$
$B$ Total available bandwidth
$b_{k}^{t}$ Bandwidth allocation ratio for client $k$ at iteration $t$
$h_{k}^{t}$ Channel gain for client $k$ at iteration $t$
$D^{t}$ Overall transmission delay at iteration $t$
$\bar{D}$ Upper bound of $D^{t}$
$\mathcal{K}^{t}$ Set of scheduled clients at iteration $t$
$m_{k}^{t}$ Error feedback memory matrix at client $k$
$\tilde{m}_{k}^{t}$ Concatenated error feedback vector
$O^{t}$ Sparsification ratio at iteration $t$
$W$ Upper bound on the singular values of LoRA
$\phi$ Constant used in the covariance bound assumption
$S$ Smoothness constant of the loss functions
$G_{k}$ Upper bound on the gradient norm at client $k$
$G$ Maximum gradient bound, $G=\max_{k}G_{k}$
$Q^{t}$ Virtual queue at iteration $t$
$D_{th}$ Average transmission delay threshold
II-A Low-rank Adaptation
Suppose we are given a pre-trained model $\theta_{P}\in\mathbb{R}^{d\times\ell}$. We aim to adapt this pre-trained model through fine-tuning for use in downstream tasks. In traditional fine-tuning, the model is initialized with the pre-trained weights $\theta_{P}$ and updated via gradient descent to obtain $\theta=\theta_{P}+\Delta\theta$, where $\Delta\theta$ represents the updates to the weights over all $d\times\ell$ parameters.
Rather than updating $\Delta\theta$ directly, LoRA seeks to represent the weight updates as a low-rank decomposition $\Delta\theta=\theta_{B}\theta_{A}$, where $\theta_{B}\in\mathbb{R}^{d\times r}$, $\theta_{A}\in\mathbb{R}^{r\times\ell}$, and $r\ll\min(d,\ell)$. In this way, the output can be represented as
$\boldsymbol{y}=\boldsymbol{x}\theta_{P}+\frac{\alpha}{r}\boldsymbol{x}\Delta\theta=\boldsymbol{x}\theta_{P}+\frac{\alpha}{r}\boldsymbol{x}\theta_{B}\theta_{A},$ (1)
where $\boldsymbol{y}\in\mathbb{R}^{\ell\times 1}$, $\boldsymbol{x}\in\mathbb{R}^{d\times 1}$, and $\frac{\alpha}{r}$ is a scaling parameter. In LoRA, the pre-trained weights $\theta_{P}$ are kept fixed during fine-tuning, and only the low-rank matrices $\theta_{A}$ and $\theta_{B}$ are updated. This significantly reduces the number of trainable parameters by a factor of $\mathcal{O}\left(\frac{r}{\min(d,\ell)}\right)$ compared to full fine-tuning.
The choice of the rank $r$ determines the fine-tuning performance, as it controls the level of approximation. A higher rank allows the model to capture more information from the original weight updates, potentially improving performance. However, it is observed that over-parameterized models often reside in a low intrinsic dimension [20, 3], meaning that there exists a maximum rank $r$ which can suffice to achieve performance comparable to fully fine-tuning.
To ensure that the weights start from the pre-trained model, $\theta_{B}$ is initialized to zero, while $\theta_{A}$ is initialized with random Gaussian values. The scaling parameter $\alpha$ controls the update magnitude, aiding in stable training. By updating only the low-rank matrices, LoRA achieves significant memory and computational efficiency, making it practical for fine-tuning large models on resource-constrained clients.
II-B Federated LoRA Fine-Tuning
Consider a network composed of a central parameter server and total $N$ local clients with single-antenna to fine-tune a global model $\theta\in\mathbb{R}^{d\times l}$ based on their local datasets using LoRA module with rank $r$. In the context of applying LoRA to FL, the pre-trained model remains fixed on both the server and the clients, so there is no need to transmit its parameters. Instead, only the LoRA $\theta_{B}$ and $\theta_{A}$ are updated locally and exchanged during each communication round. This approach significantly reduces communication overhead compared to fully fine-tuning methods that require transmitting the entire model at every round.
Each client $k\in\mathcal{N}$ owns a local dataset $\mathcal{D}_{k}=\{(x_{i},y_{i})\}_{i=1}^{D_{k}}$ with $D_{k}=|\mathcal{D}_{k}|$ data pairs such that $\mathcal{D}=\cup_{k\in\mathcal{N}}\mathcal{D}_{k}$. The local loss function of client $k$ is
$F_{k}(\theta_{B},\theta_{A})=\frac{1}{D_{k}}\sum_{x_{i}\in\mathcal{D}_{k}}f(\theta_{B},\theta_{A};x_{i},y_{i}),$ (2)
where $f(\theta_{B},\theta_{A})$ represents the sample-wise loss function. The global loss function is
$F(\theta_{B},\theta_{A})=\sum_{k\in\mathcal{K}^{t}}p_{k}F_{k}(\theta_{B},\theta_{A}),$ (3)
where $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$. To achieve a global model $\theta^{*}=\theta_{P}+\frac{\alpha}{r}\theta_{B}^{*}\theta_{A}^{*}$, clients periodically send updates to the server, assuming no direct interaction among the clients. The optimal global model $\theta^{*}$ for LoRA is identified through the following minimization process
$\theta^{*}=\underset{\theta_{B}\in\mathbb{R}^{d\times r},\theta_{A}\in\mathbb{R}^{r\times\ell}}{\arg\min}F(\theta_{B},\theta_{A}).$ (4)
Within the wireless federated LoRA fine-tuning framework, the objective function $F(\theta_{B},\theta_{A})$ is progressively minimized through a series of local and global updates.
Figure 1: System model of federated LoRA fine-tuning.
The aggregation process closely mirrors that of FedAvg, with the key difference being that only the LoRA modules are transmitted and aggregated. Specifically, the following steps are executed in the $t$-th round for $t=1,2,\dots,T$:
  1. 1.
    Broadcasting: The server distributes the $t$-th round global matrices $\theta_{B}^{t}$ and $\theta_{A}^{t}$ to the set of scheduled clients $\mathcal{K}^{t}$ by broadcasting. Since the server has enough power to transmit the global matrices correctly, we assume the downlink transmission is error-free.
  2. 2.
    Local Fine-Tuning: Each client $k\in\mathcal{K}^{t}$ initializes the global matrices as an initial point of local matrices, i.e. $\theta_{B,k}^{t}=\theta_{B}^{t},\theta_{A,k}^{t}=\theta_{A}^{t}$. Each client performs local training through gradient descent or ADAM resulting in an updated local models $\theta_{B,k}^{t+1}$ and $\theta_{A,k}^{t+1}$.
  3. 3.
    Matrices Aggregation: Client $k\in\mathcal{K}^{t}$ returns the updated local matrices $\theta_{B,k}^{t+1}$ and $\theta_{A,k}^{t+1}$ to the server. The server aggregates and averages the updates to form the $t+1$-th round global matrices as follows:
    $\theta_{B}^{t+1}=\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{B,k}^{t+1},\quad\theta_{A}^{t+1}=\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{A,k}^{t+1},$ (5)
where $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$ denotes the proportion of the dataset associated with client $k$, i.e. $\sum_{k\in\mathcal{N}}p_{k}=1$. After a total of $T$ rounds of this process, the training completes, yielding an optimally trained global model for LoRA. Although a client may intend to send the complete matrix product $\theta_{B,k}\theta_{A,k}$, to avoid the computational burden of matrix multiplication at the resource-constrained clients, it instead sends $\theta_{B,k}$ and $\theta_{A,k}$. This is because transmitting the full product would involve sending a matrix of size $d\times\ell$, which is generally much larger than $\theta_{B,k}$ and $\theta_{A,k}$ due to $r\ll\min\{d,\ell\}$. Therefore sending the two low-rank matrices not only minimizes local computation but also reduces the overall communication latency and overhead, as the total number of transmitted parameters is significantly low. Fig. 1 provides an overall framework of the wireless federated LoRA fine-tuning.
II-C Communication Model
We consider a wireless communication model between local clients and the server. In practice, the uplink, which is used by clients to transmit their LoRA matrices to the server, typically experiences more stringent throughput limitations than the downlink. Therefore, our primary focus is on the uplink communication. In contrast, the downlink benefits from higher transmit power, and thus we assume it provides reliable communication at effectively unlimited rates [5]. For the uplink communication, we consider a wireless multiple access system, specifically a frequency division multiple access (FDMA) scheme for local model transmission, with a total available bandwidth of $B$. The channel gain between the server and client is modeled as $h_{k}^{t}=s_{k}^{t}q_{k}^{-\gamma}$, where $s_{k}^{t}\sim\mathcal{N}(0,1)$ represents the small-scale channel fading coefficient, and $q_{k}^{-\gamma}$ denotes the distance-dependent path loss with exponent $\gamma$. Additionally, $b_{k}^{t}\in[0,1]$ represents the bandwidth allocation ratio for client $k\in\mathcal{N}$ where $\sum_{k=1}^{N}b_{k}^{t}=1$. The variable $a_{k}^{t}\in\{0,1\}$ indicates whether client $k$ is selected in iteration $t$, where $\sum_{k=1}^{N}a_{k}^{t}=K^{t}$ and $a_{k}^{t}=0$ implies no bandwidth is allocated to client $k$, resulting in $b_{k}^{t}=0$. The transmission rate for client $k$ can be expressed as $r_{k}^{t}=b_{k}^{t}B\log_{2}\left(1+\frac{|h_{k}^{t}|^{2}}{\sigma^{2}}\right)$, where $\sigma^{2}$ represents the variance of the additive white Gaussian noise (AWGN). During the uplink phase, we assume that the amount of information per model parameter is denoted by $v$. Thus, the communication delay for client $k$ at iteration $t$ can be written as
$D_{k}^{t}=\frac{a_{k}^{t}vr(d+\ell)}{r_{k}^{t}}=\frac{a_{k}^{t}vr(d+\ell)}{b_{k}^{t}B\log_{2}\left(1+\frac{|h_{k}^{t}|^{2}}{\sigma^{2}}\right)}.$ (6)
Note that if $a_{k}^{t}=0$, then $D_{k}^{t}=0$. Since the overall transmission delay at iteration $t$ is determined by the slowest client, it is given by $D^{t}=\max_{k\in\mathcal{K}^{t}}D_{k}^{t}.$
III Sparsified Orthogonal Fine-Tuning
III-A Loss Function Design
To reduce communication overhead in FL, model updates are commonly sparsified before transmission. An effective sparsification rule should therefore preserve the most informative components of the update. Among rank-$r$ representations of a matrix, truncated SVD provides a desirable approximation in the Frobenius norm [21], suggesting that dominant singular directions offer a principled criterion for importance. However, computing SVD at every communication round is prohibitively expensive for resource-constrained clients in LoRA-based FL. We therefore seek a lightweight surrogate that captures this structural property without explicit decomposition.
A fundamental component of our sparsification method is the design of a loss function that enforces nearly orthogonal properties on the matrices $\theta_{B,k}$ and $\theta_{A,k}$. Specifically, we aim for $\theta_{B,k}^{\textsf{T}}\theta_{B,k}$ and $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ to approximate diagonal matrices, analogous to the roles of $U$ and $V$ in SVD. Achieving this ensures that the column vectors of $\theta_{B,k}$ and the row vectors of $\theta_{A,k}$ are nearly orthogonal, thereby simplifying subsequent computations. To accomplish this, each client $k$ constructs the following loss function:
$L_{k}=L_{task,k}$ $+\zeta(\|\theta_{B,k}^{\textsf{T}}\theta_{B,k}-diag(\theta_{B,k}^{\textsf{T}}\theta_{B,k})\|_{F}^{2}$
$+\|\theta_{A,k}\theta_{A,k}^{\textsf{T}}-diag(\theta_{A,k}\theta_{A,k}^{\textsf{T}})\|_{F}^{2}),$ (7)
where $L_{task,k}$ denotes the task-specific loss for client $k$, $\zeta$ is a scaling parameter, and $\text{diag}(C)$ denotes a diagonal matrix with the diagonal elements of $C$. This loss function ensures that $\theta_{B,k}$ and $\theta_{A,k}$ approximate the orthogonal properties of the $U$ and $V$ matrices in SVD, respectively. The effect of the proposed loss function and the resulting performance are examined in detail in Subsection V-D. Therefore, this property allows us to compute the Frobenius norm by aggregating the $\ell_{2}$-norms of the individual vectors across the $r$ components, as expressed mathematically
$\|\theta_{B,k}\theta_{A,k}\|_{F}^{2}=\sum_{i=1}^{r}\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2},$ (8)
where $\theta_{B,k}[:,i]\!$ and $\theta_{A,k}[i,:]\!$ are the $i$-th column and row vectors, respectively, whose near orthogonality ensures that cross terms vanish, simplifying the calculation to sum of vector norms.
III-B Sparsification Method
At the beginning of the iteration, each client $k$ performs local training using the proposed loss function (III-A), resulting in $\theta_{B,k}$ and $\theta_{A,k}$. The client then applies sparsification to updated matrices. Specifically, the client sets the magnitude of all elements in $\theta_{B,k}$ and $\theta_{A,k}$ to zero, except for the selected $r(d+\ell)O_{k}$ elements, where $O_{k}\in[0,1]$ is the sparsification ratio of client $k$.
Since $\theta_{B,k}$ and $\theta_{A,k}$ exhibit nearly orthogonal properties, we can compute singular values by calculating the norms of the matrices $\theta_{B,k}$ and $\theta_{A,k}$. For instance, the $i$-th singular value, where $i=\{1,...,r\}$, can be obtained as $\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2}$. Because the magnitude of singular values represents the importance of their corresponding singular vectors, we implement sparsification based on the magnitude of these singular values. For each singular vector, we define $o_{k,i}$ as the number of non-zero elements retained in both $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$. The value of $o_{k,i}$ is determined by distributing the total number of non-sparse elements proportionally to the magnitude of the singular values. Specifically, $o_{k,i}$ is calculated as follows:
$o_{k,i}=\frac{O_{k}r(d+\ell)\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2}}{\sum_{i=1}^{r}\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2}}.$ (9)
Figure 2: Computational complexity comparison between SOFT and the conventional full-multiplication and SVD-based approach. (a) Complexity versus the LoRA rank $r$. (b) Complexity versus matrix dimension.
Finally, based on (9), client $k$ selects top-$o_{k,i}$ elements in $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$. Although the proposed sparsification method also requires calculating the ll-2 norm $r$ times in each iteration, this computation has a complexity of $\mathcal{O}(r(d+\ell))$, negligible compared to the original complexity caused by full matrix multiplication and SVD operation, i.e. $\mathcal{O}(dr\ell+\min(d,\ell)\cdot d\ell)$. To further illustrate the computational efficiency of SOFT, Fig. 2 compares its computational complexity with that of the conventional full-multiplication and SVD-based approach under varying LoRA ranks and matrix dimensions. In Fig. 2, we fix the matrix dimensions as $d=\ell=4096$ and vary the LoRA rank $r$, whereas in Fig. 2, we fix the LoRA rank as $r=8$ and vary the matrix dimension with $d=\ell$. The $y$-axis is plotted on a logarithmic scale due to the large gap in complexity between the two approaches. This result confirms that the additional overhead of the proposed sparsification method is negligible compared to the original computation.
Remark 1.
Since the loss function (III-A) enforces $\theta_{B}^{\textsf{T}}\theta_{B}$ and $\theta_{A}\theta_{A}^{\textsf{T}}$ to be approximately diagonal, the cross terms in $\|\theta_{B}\theta_{A}\|_{F}^{2}$ vanish and it holds $\|\theta_{B}\theta_{A}\|_{F}^{2}=\sum_{i=1}^{r}\|\theta_{B}[:,i]\|_{2}^{2}\|\theta_{A}[i,:]\|_{2}^{2},$ so the per-rank score $\|\theta_{B}[:,i]\|_{2}^{2}\|\theta_{A}[i,:]\|_{2}^{2}$ acts as a surrogate singular value. Performing the sparsification with the per-rank non-zeros $o_{k,i}\propto\|\theta_{B}[:,i]\|_{2}^{2}\|\theta_{A}[i,:]\|_{2}^{2}$ is therefore a consistent SVD-proxy selection rule. Although the loss function (III-A) can slightly reduce accuracy when used alone by encouraging near orthogonality, incorporating the subsequent sparsification step improves overall performance and thus produces favorable accuracy–efficiency trade-offs.
III-C Error Feedback with SOFT
While the proposed sparsification method effectively reduces communication costs by transmitting only the most significant elements of $\theta_{B,k}$ and $\theta_{A,k}$, it inevitably introduces sparsification errors. To mitigate the impact of these errors on model convergence and performance, we incorporate an error feedback mechanism into SOFT. The error feedback mechanism works by accumulating the sparsification errors at each client and incorporating them into the updates [8]. Specifically, each client maintains local error memory matrices, denoted as $m_{B,k}$ and $m_{A,k}$, which store the cumulative sparsification errors for $\theta_{B,k}$ and $\theta_{A,k}$, respectively. The sparsified update can be expressed as
$\theta_{B,k}^{t}=\mathcal{S}(m_{B,k}^{t}+\Delta\theta_{B,k}^{t}),\theta_{A,k}^{t}=\mathcal{S}(m_{A,k}^{t}+\Delta\theta_{A,k}^{t}),$ (10)
where $\mathcal{S}$ operates the sparsification algorithm, $\Delta\theta_{B,k}^{t}$ and $\Delta\theta_{A,k}^{t}$ represent obtained model of client $k$ at time $t$ using the loss function (III-A). Subsequently, the error memory is updated to accumulate the residuals resulting from sparsification:
$m_{B,k}^{t+1}=m_{B,k}^{t}+\Delta\theta_{B,k}^{t}-\theta_{B,k}^{t},$ (11)
$m_{A,k}^{t+1}=m_{A,k}^{t}+\Delta\theta_{A,k}^{t}-\theta_{A,k}^{t}.$ (12)
For clarity and brevity, we define the concatenated representations as follows:
$\tilde{m}_{k}^{t}=\begin{bmatrix}{m_{B,k}^{t\textsf{T}}}m_{A,k}^{t}\end{bmatrix}$,
$\Delta\tilde{\theta}_{k}^{t}=\begin{bmatrix}{\Delta\theta_{B,k}^{t\textsf{T}}}\Delta\theta_{A,k}^{t}\end{bmatrix},\quad\tilde{\theta}_{k}^{t}=\begin{bmatrix}{\theta_{B,k}^{t\textsf{T}}}\theta_{A,k}^{t}\end{bmatrix}$, (13)
where $\cdot$] is concatenation operation. With these definitions, the overall error memory update can be succinctly written as
$\tilde{m}_{k}^{t+1}=\tilde{m}_{k}^{t}+\Delta\tilde{\theta}_{k}^{t}-\tilde{\theta}_{k}^{t}.$ (14)
This process effectively corrects the errors introduced by sparsification, ensuring that the model updates are more accurate over time. The error feedback mechanism enhances the convergence of the FL process despite the sparsification. The advantage of SOFT will be evaluated in Subsection V-A.
We now state a lemma that shows the errors caused by the SOFT using the following assumption[8].
Assumption 1.
Let $\mathcal{S}:\mathbb{R}^{q}\rightarrow\mathbb{R}^{q}$ be a sparsification operator and $u$ be the number of non-zero elements satisfying $0<u<q.$ Then, for all $x\in\mathbb{R}^{q}$, the operator $\mathcal{S}$ satisfies $\mathbb{E}\left[\|\mathcal{S}(x)-x\|_{2}^{2}\right]\leq\left(1-\frac{u}{q}\right)\|x\|_{2}^{2}.$
Lemma 1.
Under nearly orthogonal properties, sparsification rate $O_{k}$ and Assumption 1, the expected squared norm of the sparsification error $\tilde{m}_{k}^{t+1}$ at iteration $t$ is upper-bounded as
$\mathbb{E}\left[\left\|\tilde{m}_{k}^{t+1}\right\|_{F}^{2}\right]\leq\frac{4(1-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}.$ (15)
Proof:
See Appendix A. ∎
Corollary 1.
Under the same assumptions and conditions as in Lemma 1, the sparsification error $m_{k}^{t+1}\in\mathbb{R}^{d\times\ell}$, i.e. $m_{k}^{t+1}=m_{B,k}^{t}m_{A,k}^{t}$ at iteration $t$ is upper-bounded as
$\mathbb{E}\left[\left\|m_{k}^{t+1}\right\|_{F}\right]\leq\frac{2(1-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}.$ (16)
Proof:
Applying the Arithmetic and Geometric Mean inequality (AM-GM inequality), it follows that
$2\|m_{B,k}^{t+1}m_{A,k}^{t+1}\|_{F}\leq\|m_{B,k}^{t+1}\|_{F}^{2}\!+\!\|m_{A,k}^{t+1}\|_{F}^{2}\!=\!\|\tilde{m}_{k}^{t+1}\|_{F}^{2}.$ (17)
Invoking Lemma 1, we have
$\mathbb{E}\left[\left\|m_{k}^{t+1}\right\|_{F}\right]\leq\frac{2(1-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}.$ (18)
IV Two-stage Federated Fine-Tuning
IV-A Convergence Analysis and Problem Formulation
In LoRA-based FL, unlike conventional FL approaches, the use of LoRA modules introduces two additional factors that must be considered in the learning process. First, as discussed in Subsection II-A, the performance of LoRA is predominantly influenced by the rank $r$. Therefore, an analysis of the impact of the selected rank for fine-tuning should be included. To relate the fully fine-tuned update to the rank $r$ LoRA update, following [22], we introduce the following assumption in the form of an inequality.
Assumption 2.
Let $\theta_{o}$ denote the LoRA model that achieves the same performance as the fully fine-tuned model, and let $\theta_{r}$ be the LoRA model of rank $r$ trained using the loss function (III-A). Denote $\sigma(\theta_{r})$ as a proxy for the dominant scale of the rank-$r$ LoRA update, i.e. $\sigma(\theta_{r})=\max_{i}\mathbb{E}\|\theta_{B}[:,i]\|_{2}\|\theta_{A}[i,:]\|_{2}$, for $i=1,\dots,r$. Then, there exists a constant $H>0$ such that the LoRA error satisfies:
$\mathbb{E}[\|\theta_{o}-\theta_{r}\|_{F}^{2}]\leq H(r_{\max}-r)\sigma^{2}(\theta_{r}),$ (19)
where $r_{\max}$ is the rank of $\theta_{o}$.
Remark 2.
Assumption 2 is inspired by the SVD. Since $\theta_{r}$ is trained to maximize performance for a given rank $r$, it is expected that the dominant singular components of $\theta_{o}$ are well approximated by $\theta_{r}$. In other words, the first $r$ singular values of $\theta_{r}$ are assumed to closely match those of $\theta_{o}$. Therefore, $\|\theta_{o}-\theta_{r}\|_{F}^{2}$ can be upper bounded by $(r_{\max}-r)$ copies of the squared $r$-th largest singular value. This behavior indicates that the error between the optimal model $\theta_{o}$ and the rank $r$ LoRA model $\theta_{r}$ diminishes with increasing $r$.
Second, as mentioned in Subsection II-B, clients transmit the matrices $\theta_{B}$ and $\theta_{A}$ instead of the full matrix product $\theta_{B}\theta_{A}$. While this approach reduces computational overhead, it introduces an unexpected additional term when aggregating the updates at the server [17, 23]. Specifically, since the server averages $\theta_{A,k}$ and $\theta_{B,k}$ independently, it introduce mathematical errors to the global model, i.e., $\mathbb{E}_{k}[\theta_{B,k}\theta_{A,k}]\neq\mathbb{E}_{k}[\theta_{B,k}]\mathbb{E}_{k}[\theta_{A,k}]$. The additional noise term introduced in the aggregation at the server can be expressed by the covariance term $Cov(\cdot)$ as follows.
$\mathbb{E}_{k}$ $[\theta_{B,k}\theta_{A,k}]-\theta_{B}\theta_{A}$
$=\mathbb{E}_{k}[(\theta_{B,k}-\theta_{B})(\theta_{A,k}-\theta_{A})]$ (20)
$=\sum_{i=1}^{r}\mathbb{E}\left[\left(\theta_{B,k}[:,i]-\theta_{B}[:,i]\right)\left(\theta_{A,k}[i,:]-\theta_{A}[i,:]\right)\right]$ (21)
$=\textit{Cov}(\theta_{B},\theta_{A}).$ (22)
As observed in (IV-A), $\textit{Cov}(\theta_{B},\theta_{A})$ reflects the discrepancy between the aggregated model and the individual client models. In FL, each client begins parameter refinement from a common initial model. Consequently, the discrepancy increases when the data distributions across clients are more heterogeneous [24], since the locally updated models tend to diverge in different directions, leading to an increase in $\textit{Cov}(\theta_{B},\theta_{A})$. Furthermore, the model size plays a critical role in this phenomenon. Larger models possess a higher number of parameters, which naturally leads to a greater overall parameter magnitude measured by the Frobenius norm $\|\theta_{r}\|_{F}^{2}$.
To formally account for this behavior, we introduce the following proposition:
Assumption 3.
Let $\theta_{r}$ denote the LoRA model with rank $r$. The expected Frobenius norm of the covariance of $\theta_{B}$ and $\theta_{A}$ satisfies $\mathbb{E}[\|\textit{Cov}(\theta_{B},\theta_{A})\|_{F}^{2}]\leq\phi\|\theta_{r}\|_{F}^{2},$ where $\phi$ is a constant that quantifies the degree of data heterogeneity across the clients.
Remark 3.
Assumption 3 indicates that, in the LoRA-based FL framework, the covariance between $\theta_{B}$ and $\theta_{A}$ is significantly affected by both the diversity of client data and the size of the model. Empirical evidence supporting this relationship is provided in Subsection V-D.
Next, for theoretical analysis, we make the following assumptions typically made in analyzing FL family [25].
Assumption 4.
The local objective function $F_{k}(\theta)$ is $S$-smooth, i.e. $\|\nabla F_{k}(\theta)-\nabla F_{k}(\theta^{\prime})\|_{F}\leq S\|\theta-\theta^{\prime}\|_{F},$ for all $\theta$. It also satisfies $F_{k}(\theta^{\prime})\leq F_{k}(\theta)+\nabla F_{k}(\theta)^{\textsf{T}}(\theta^{\prime}-\theta)+\frac{S}{2}\|\theta^{\prime}-\theta\|_{F}^{2},$ for all $\theta$. The global objective function $F(\theta)$, being the average of the local objectives, is then also $S$-smooth and the global objective function $F(\theta)$ is lower bounded such as $F(\theta)\geq F(\theta^{*})$ for all $\theta$.
Assumption 5.
The stochastic gradient at each client is unbiased such as $\mathbb{E}[\nabla F_{k}^{\scriptstyle{t,e}}(\theta)]=\nabla F_{k}(\theta)$ for all $k,t,e$.
Assumption 6.
The expected squared norm of the stochastic gradient and the largest singular value of LoRA model at each client are upper bounded as $\mathbb{E}[\|\nabla F_{k}^{\scriptstyle(t,e)}(\theta)\|_{F}^{2}]\leq G_{k}^{2}$ and $\sigma(\theta_{r})\leq W$, respectively. Hence, the Frobenius norm of the LoRA weight is bounded as $\mathbb{E}[\|\theta_{r}\|_{F}^{2}]\leq rW^{2}$.
Theorem 1.
Based on Assumptions 2, 3, 4, 5, 6 and Lemmas 1 and 3, the optimality gap of LoRA-based FL after $T$ global iterations is upper-bounded as
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\|\nabla F_{r}(\theta_{r}^{t})\right\|_{F}^{2}\right]\leq\frac{2}{\eta T}\left(\mathbb{E}\left[F_{r}(\theta_{r}^{0})-F^{*}\right]\right)$
$+\frac{1}{T}\sum_{t=0}^{T-1}\bigg(2S^{2}H(r_{\max}-r)W^{2}+4S^{2}\phi rW^{2}+\eta SE^{2}G^{2}$
$+\frac{8(N-K^{t})}{K^{t}(N-1)}\eta^{2}S^{2}E^{2}G^{2}+\frac{4N(1-O^{t})^{2}}{K^{t}{(O^{t})}^{4}}rS^{2}W^{4}$
$+\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}G^{2}\bigg)=\gamma^{T},$ (23)
where $G^{2}=\max_{k}G_{k}^{2}$; $E$ and $\eta$ represent the number of local iterations and learning rate, respectively.
Proof:
See Appendix B. ∎
Remark 4.
Theorem 1 indicates that the convergence behavior of the LoRA-based FL is governed by several interconnected parameters: the sparsification ratios $O^{t}$ and the LoRA rank $r$. These factors are intrinsically linked, collectively influencing the trade-off between communication efficiency and learning performance. For instance, a higher sparsification ratio $O^{t}$ reduces communication overhead but may degrade convergence, while increasing the rank $r$ enhances model expressiveness at the cost of additional communication. By carefully balancing these factors, we can enhance both communication efficiency and learning effectiveness within the constraints of the LoRA-based FL environment.
Building upon the convergence analysis presented in Theorem 1, our goal is to jointly optimize the key parameters to minimize the convergence upper bound. Specifically, we aim to determine the optimal LoRA rank $r$, sparsification ratios $O^{t}$, and bandwidth allocations $b_{k}$ to enhance both communication efficiency and learning performance under the constraints of limited bandwidth and training time. Our optimization problem is formulated as follows.
$\mathcal{P}_{1}$ $\min_{r,O^{t},b_{k}^{t}}\gamma^{T},$ (24a)
s.t. $\ \ r\leq r_{\max},r\in\mathbb{Z}^{+}$ (24b)
$\ \ \frac{1}{T}\sum\nolimits_{t=0}^{T-1}D^{t}\leq D_{th},$ (24c)
$\ \ O_{min}<O_{k}^{t}\leq 1,$ (24d)
$\ \ 0\leq b_{k}^{t}\leq 1,\quad\sum\nolimits_{k=1}^{N}b_{k}^{t}=1,$ (24e)
where (24b) is the rank feasibility constraint; (24c) is the average transmission delay constraint  [^1] In this paper, the delay constraint denotes the latency to upload a client’s model to the server in each iteration.; (24d) is the sparsification ratio constraint with minimum sparsification ratio $O_{min}$; (24e) is the constraints for the bandwidth allocation strategy.
Problem $\mathcal{P}_{1}$ poses two main challenges. First, the structural LoRA rank $r$ must be determined before training because it fixes the model structure and directly affects the model performance as well as the communication payload. Therefore, unlike $O_{k}^{t}$ and $b_{k}^{t}$, $r$ is not an online control variable. Although some prior works [16, 17, 18] consider adaptive rank strategies, they still operate within a pre-defined structural rank budget and do not resolve the structural rank selection problem considered here. Moreover, selecting $r$ online in wireless FL would require repeated evaluation of candidate ranks under time-varying CSI, leading to prohibitive training and signaling overhead. Second, $\mathcal{P}_{1}$ is a MINLP with long-term latency constraints and strongly coupled variables, which makes direct online optimization intractable in practice. We therefore adopt a two-stage design: $r$ is selected offline before training, while $O_{k}^{t}$ and $b_{k}^{t}$ are optimized online during training.
IV-B Offline Stage
As discussed in the previous subsection, the offline stage is conducted prior to training with the specific goal of optimizing the appropriate LoRA rank $r$. Since the actual CSI cannot be obtained during the offline stage, the channel is approximated by assuming the average channel as the actual channel. The initial sparsification ratio is set to $O^{0}$. The offline stage thus considers the following approximated problem:
$\mathcal{P}_{2}$ $\min_{r,O^{0},b_{k}^{0}}\gamma^{0},$ (25a)
s.t. $\ \ r\leq r_{\max},~r\in\mathbb{Z}^{+}$ (25b)
$\ \ D^{0}\leq D_{th},$ (25c)
$\ \ O_{min}<O^{0}\leq 1,$ (25d)
$\ \ 0\leq b_{k}^{0}\leq 1,\quad\sum\nolimits_{k=1}^{N}b_{k}^{0}=1,$ (25e)
where $\gamma^{0}$ is the upper-bound on the initial optimality gap in Theorem 1.
Note that the Problem $\mathcal{P}_{2}$ is still an NP-hard MINLP problem due to the discrete rank value and non-convexity caused by the coupled variables. To address these challenges, we first fix the integer variable $r$ to a specific value and solve the resulting continuous optimization problem $O^{0}$. By iterating over all feasible values of $r$ within the constraint, we can identify the optimal integer $r$. To efficiently solve the continuous optimization problem for a fixed $r$, we present the following lemma, which provides the optimal values of $O^{0}$ given an integer $r$.
Lemma 2.
In Problem $\mathcal{P}_{2}$, given an integer value $r$, $O^{0}$ that minimize the objective function can be obtained as
$O^{0}=\max\left(O_{min},\min\left(1,\frac{ND_{th}}{AK^{0}r}\right)\right),$ (26)
where $A=\sum_{k=1}^{N}\frac{v(d+\ell)}{B\log_{2}(1+|h_{k}|^{2}/\sigma^{2})}$.
Proof:
[…]
$D_{th}$ $=\mathbb{E}\left[\sum_{k\in\mathcal{K}^{0}}\frac{vO^{0}r(d+\ell)}{B\log_{2}(1+\frac{|h_{k}|^{2}}{\sigma^{2}})}\right]$ (27)
$=\frac{K^{0}}{N}\sum_{k=1}^{N}\frac{vO^{0}r(d+\ell)}{B\log_{2}(1+\frac{|h_{k}|^{2}}{\sigma^{2}})}.$ (28)
Consequently, using $O_{min}\!<O^{0}\leq 1$, we can obtain
$O^{0}=\max\left(O_{min},\min\left(1,\frac{ND_{th}}{AK^{0}r}\right)\right),$ (29)
where $A=\sum_{k=1}^{N}\frac{v(d+\ell)}{B\log_{2}(1+|h_{k}|^{2}/\sigma^{2})}$. ∎
Based on Lemma 2, we can obtain the optimal LoRA rank to be used in training. Subsequently, clients perform fine-tuning based on the specified LoRA rank $r$. At each iteration, an online-stage optimization is conducted to determine the values of $O^{t}$ and $b_{k}^{t}$.
IV-C Online Stage
In this subsection, we address the optimization of the remaining parameters for the specified LoRA rank. Our objective is to dynamically adjust the sparsification ratio $O^{t}$, and the bandwidth allocation $b_{k}^{t}$} for every iteration $t$. These adjustments are made in response to real-time network conditions and the availability of resources. Since these parameters are interdependent and subject to long-term latency constraints, Lyapunov optimization is employed to perform online optimization at each iteration. To handle the time-average latency constraint specified in (24c), we introduce a virtual queue $Q^{t}$ that evolves according to the following update rule
$Q^{t+1}=\max(Q^{t}+D^{t}-D_{th},0),$ (30)
where $Q^{0}=0$. The virtual queue $Q^{t}$ effectively tracks the accumulation of latency violations over time. Ensuring the stability of this queue is equivalent to satisfying the long-term latency constraint. To measure the congestion of queue $Q^{t}$, the Lyapunov function is defined as
$L(Q^{t})=\frac{1}{2}(Q^{t})^{2}.$ (31)
To bound the increase of the virtual queues, we need to constrain the expected increase of the Lyapunov function. Thus the Lyapunov drift at iteration $t$ is
$\Delta(Q^{t})=\mathbb{E}[L(Q^{t+1})-L(Q^{t})|Q^{t}],$ (32)
where the expectation is taken over the random system state, i.e. channel state. To balance queue stability with the minimization of the convergence upper bound $\gamma^{t}$, we introduce the drift-plus-penalty function
$\Delta(Q^{t})+V\mathbb{E}\left[\gamma^{t}|Q^{t}\right],$ (33)
where $V>0$ is a control parameter that weights the importance of the convergence performance relative to queue stability. A larger $V$ places more emphasis on optimizing the convergence bound. To make the optimization tractable, we derive an upper bound for the Lyapunov drift. Expanding the drift expression, we have:
$\Delta(Q^{t})$
$=\mathbb{E}\left[\frac{1}{2}(Q^{t+1})^{2}-\frac{1}{2}(Q^{t})^{2}\Big|Q^{t}\right]$ (34)
$=\mathbb{E}\left[\frac{1}{2}\left(\max\left\{Q^{t}+D^{t}-D_{th},0\right\}\right)^{2}-\frac{1}{2}(Q^{t})^{2}\Big|Q^{t}\right]$ (35)
$\leq\mathbb{E}\left[\frac{1}{2}\left(Q^{t}+D^{t}-D_{th}\right)^{2}-\frac{1}{2}(Q^{t})^{2}\Big|Q^{t}\right]$ (36)
$=Q^{t}\mathbb{E}\left[D^{t}-D_{th}|Q^{t}\right]+\frac{1}{2}\mathbb{E}\left[\left(D^{t}-D_{th}\right)^{2}|Q^{t}\right]$ (37)
$\leq B+Q^{t}\left(\mathbb{E}\left[D^{t}|Q^{t}\right]-D_{th}\right),$ (38)
where $B=\dfrac{1}{2}(\bar{D}-D_{th})^{2}$, and $\bar{D}$ represents the upper bound of all possible $D^{t}$. Note that $\bar{D}$ can be empirically determined, for example, by assuming that all clients receive the minimum possible bandwidth and experience the worst-case channel conditions. The inequality in (36) holds because $(\max{x,0})^{2}\leq x^{2},;\forall x\in\mathbb{R}$, while the inequality in (38) follows from the fact that $D^{t}\leq\bar{D}$.
By substituting the upper bound Lyapunov drift into the drift-plus-penalty function, we obtain an expression that we can optimize at each time slot. Note that since the current queue $Q^{t}$ is already observed at the beginning of time slot $t$, the expectation in the drift is over $D^{t}$ which affects the transition $Q^{t}$ to $Q^{t+1}$. Also, under the assumption of perfect knowledge of CSI at the server, the channel randomness at time $t$ is fully observed at the beginning of the time slot. Consequently, both $D^{t}$ and $\gamma^{t}$ become deterministic functions of the observed CSI at time $t$. Discarding constant terms that do not affect the optimization, we reformulate the per-slot problem based on the drift-plus-penalty function as follows:
$\mathcal{P}_{3}:$ $\min_{O^{t},b_{k}^{t}}Q^{t}D^{t}+V\gamma^{t}$ (39a)
s.t. $\ \ O_{\min}\leq O^{t}\leq 1,$ (39b)
$\ \ 0\leq b_{k}^{t}\leq 1,\quad\sum\nolimits_{k=1}^{N}b_{k}^{t}=1,$ (39c)
where $\gamma^{t}$ is the upper-bound on the optimality gap at time $t$ in Theorem 1.
To address Problem $\mathcal{P}_{3}$, we begin by demonstrating that it is convex with respect to $O^{t}$. Specifically, the second derivative of the objective function is given by
$\frac{8NrS^{2}W^{2}}{K^{t}{O^{t}}^{6}}(3{O^{t}}^{2}-12O^{t}+10),$ (40)
which is always positive due to the constraint $O_{\text{min}}<O^{t}\leq 1$. This confirms the convexity of the problem, allowing us to employ efficient convex optimization algorithms or well-established software tools such as CVX to find the global optimum. Once $O^{t}$ is determined, the overall delay $D^{t}$ for the current iteration is set. Based on (6) and $D^{t}=\max_{k\in\mathcal{K}^{t}}D^{t}_{k}$, we can now determine the bandwidth allocation strategy $b_{k}^{t}$. We leverage the property that only selected clients, i.e. those with $a_{k}^{t}=1$, receive nonzero bandwidth, i.e. $b_{k}^{t}>0$, which allows us to derive
$b_{k}^{t}=\frac{vO^{t}r(d+\ell)}{BD^{t}\log_{2}\left(1+\frac{|h_{k}|^{2}}{\sigma^{2}}\right)}=\frac{1}{A^{\prime}\log_{2}\left(1+\frac{|h_{k}|^{2}}{\sigma^{2}}\right)},$ (41)
where $A^{\prime}=\sum_{k\in\mathcal{K}^{t}}\frac{1}{\log_{2}\left(1+\frac{|h_{k}|^{2}}{\sigma^{2}}\right)}$. This expression indicates that the bandwidth allocation for each client is inversely proportional to the logarithm of their channel gain-to-noise ratio, promoting fairness among clients with varying channel conditions.
At each iteration, the server computes the optimal values of $b_{k}^{t}$ and $O^{t}$ and communicates these parameters to the selected clients. The clients then utilize the allocated bandwidth and parameters for data transmission.
V Numerical Results
In this section, we present simulation results to demonstrate the effectiveness of the proposed SOFT and TSFA framework. All experiments were conducted using Python 3.8 on an Ubuntu server equipped with NVIDIA GeForce RTX 3090 GPUs. The total number of clients is 100, and 10 clients are uniformly sampled in each round in all experiments.
(a) LoRA rank $r=4$
(b) LoRA rank $r=8$
(c) OSFA and TSFA
Figure 3: Test accuracy of CIFAR-100 classification. (a) and (b) show the performance using LoRA with ranks $r=4,r=8$, respectively, under various sparsification methods, while panel (c) compares the performance of OSFA and TSFA.
We evaluated the performance of our method on the CIFAR-100 datasets [26]. In our experiments, we utilized transformer-based pre-trained models for fine-tuning. Specifically, we employed the pre-trained ViT-Base model, which stands for the Vision Transformer Base model. ViT is a transformer-based architecture specifically designed for image recognition tasks. LoRA layers are inserted into Blocks 0, 3, 6, and 9 of the ViT-Base model. Training was performed using the AdamW optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-4}$. The local batch size and the number of local epochs were set to 8 and 1, respectively. Regarding the baseline schemes in the simulation, we consider the following frameworks:
  • FLASC[19]: Each client selects and transmits only the elements of the LoRA matrices with the largest magnitudes. Specifically, after computing the low-rank adaptation matrices, clients identify the top-$q$ elements with the highest absolute values and sparsify the rest.
  • RLoRA[8]: Each client selects and transmits only the elements of the LoRA matrices randomly without considering their magnitudes or positions within the matrices.
  • HetLoRA[16]: Each client selects and transmits only the elements of the LoRA matrices in a structured manner. Specifically, each client preserves the elements corresponding to lower rank indices and applies sparsification to those associated with higher rank indices. Here, rank indice refers to the ordering of LoRA components as defined by the model’s architecture.
  • FedLoDrop[15]: Each client employs a structured random sparsification strategy on the LoRA matrices. Specifically, the same randomly selected LoRA rank components are retained in both matrices, so that the corresponding columns of the left LoRA matrix and rows of the right LoRA matrix are transmitted while the others are masked to zero. Unlike the original FedLoDrop, which uses Bernoulli sparsification and results in a stochastic transmission size, we enforce an exact communication budget by retaining precisely the number of rank components specified by the target sparsification ratio.
  • ILoRA[3]: The Ideal LoRA framework represents an idealized environment where there are no communication constraints between clients and the server. In this scenario, clients transmit full LoRA matrices without any sparsification. This serves as an upper bound for performance.
V-A Effects of SOFT
(a) Task loss only
(b) Proposed loss
(c) Test accuracy comparison between (a) and (b)
Figure 4: $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ of Vit-Base blocks.6.mlp.fc1 layer: (a) without and (b) with the proposed loss function and (c) comparision of their test accuracy.
In this subsection, we evaluate the effectiveness of the SOFT method by comparing it with the baseline methods. Figs. 3(a) and 3(b) illustrate the test accuracy on the CIFAR-100 dataset using various sparsification methods with LoRA rank $r=4$ and $r=8$, respectively. We set the sparsification ratio as $0.5$ for all baselines and the proposed schemes to ensure a fair comparison under the same communication budget. Notably, SOFT consistently maintains higher accuracy compared to the other schemes. This superior performance can be attributed to SOFT’s ability to dynamically sparsify the LoRA modules by considering the importance of individual parameters. In LoRA module-based fine-tuning, two matrices are multiplied to form a single layer. FLASC exhibits lower performance in this context. This is because simply selecting the top-$q$ elements does not effectively capture the crucial interactions between the two matrices involved in LoRA. As a result, the essential information required for accurate model updates may be lost, leading to suboptimal performance.
Similarly, HetLoRA, which transmits only the elements associated with the lower rank indices throughout training, underperforms SOFT. Since HetLoRA sets a fixed sparsity pattern without adapting to the varying importance of parameters during training, it cannot capture the dynamic changes in the model’s weight distribution. This rigidity results in less effective learning and lower accuracy. Furthermore, RLoRA and FedLoDrop are not well suited for LoRA-based fine-tuning. RLoRA randomly selects parameters for transmission without considering their importance, which can impair the learning process and hinder convergence. Similarly, FedLoDrop randomly sparsifies LoRA rank components without accounting for their contribution to the overall low-rank update. As a result, the transmitted updates may fail to preserve the most informative components, leading to significant degradation in update quality.
V-B Effects of Offline and Online Stage
In Section IV, we present an optimization framework that employs a two-stage process: an offline stage for pre-training configuration and an online stage for dynamic parameter adjustment during training. In this subsection, we evaluate the performance of TSFA by comparing it with a one-stage variant (OSFA), which forgoes pre-optimization of the LoRA rank $r$ and instead relies solely on online adjustments with arbitrarily chosen $r$ values. The experiments were conducted using $N=100$, $K^{t}=10$, $D_{th}=500s$ and $V=0.0001$ under a Rayleigh block fading channel characterized by a zero mean and unit variance, with constant channel gains during each user uplink transmission. The SOFT was utilized for sparsification.
Fig. 3(c) clearly demonstrates that pre-selecting an appropriate LoRA rank $r$ is essential for optimal performance. When OSFA with $r=1$, $r=2$ or $r=4$, the model suffers from insufficient capacity, despite benefiting from lower sparsification ratios compared to TSFA. In these cases, the reduced rank is inadequate to capture the model’s complexity, leading to under-parameterization and poor learning of the underlying data distribution. Conversely, when OSFA employs higher $r$ values such as $r=16$, $r=32$, or $r=64$, the increased model capacity requires a corresponding increase in the sparsification ratio to satisfy communication constraints. This, in turn, results in excessive sparsification, which limits the transmission of critical parameter updates and adversely affects learning. These findings underscore the importance of the offline stage in TSFA. By determining an appropriate LoRA rank $r$ prior to training, the two-stage approach effectively balances model capacity and communication efficiency, thereby avoiding the pitfalls of both under-parameterization and over-sparsification.
V-C Effects of Proposed Loss Function
Figs. 4(a) and 4(b) illustrate the effect of our proposed loss function on promoting near orthogonality within $\theta_{B,k}$ and $\theta_{A,k}$. After training with the proposed loss function as shown in Fig. 4(b), the product $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ becomes approximately diagonal, demonstrating the successful enforcement of nearly orthogonal properties. Furthermore, Fig. 4(c) compares the accuracy of CIFAR-100 between with and without the proposed orthogonal loss. Adding the orthogonal loss results in negligible performance degradation and yields accuracy nearly identical to that of the task-loss-only baseline. These results indicate that our method enables effective sparsification without requiring full matrix multiplication or SVD operations, thereby maintaining performance while significantly reducing computational overhead.
Figure 5: Frobenius norm of covariance under different number of shards.
Figure 6: Test accuracy of CIFAR-100 under different number of shards.
V-D Effects of Covariance
In Section II, we introduced the covariance term that naturally arises due to the separate transmission and aggregation of LoRA matrices. In this subsection, we provide empirical evidence illustrating the strong correlation between this covariance term and the degree of data heterogeneity in a non-IID setting. To simulate a heterogeneous environment, we employ a shard-based non-IID partitioning strategy. We begin by sorting the dataset by class labels, ensuring that each subgroup is composed of samples belonging to the same class. These sorted subsets are then split into a predetermined number of shards, with each shard predominantly containing data from a single class. By randomly distributing these shards to clients, each client receives data biased toward certain classes, thereby introducing data heterogeneity.
Fig. 5 shows how the covariance term evolves during training for different shard sizes, and Fig. 6 shows the corresponding test accuracy. We set $r=32$ and assume no communication constraint in order to sorely consider non-IID effects. As the shard size increases, the sampling process becomes increasingly similar to an IID scenario. Larger shards yield more balanced client data distributions, which naturally reduce the discrepancy among locally updated LoRA parameters. This results in a lower covariance magnitude, as evidenced by the decreasing trend in Fig. 5. The results in Fig. 6 confirm this relationship, scenarios with lower covariance consistently achieve higher test accuracy, highlighting the pivotal role that covariance plays in model convergence and generalization under non-IID conditions. Developing methods to reduce or compensate for this effect whether through data-driven approaches, adaptive averaging mechanisms, or improved LoRA configurations, remains an intriguing avenue for future research.
Figure 7: Test accuracy of AG News dataset with rank $r=8$.
V-E Results on Other Datasets and Models
To verify that our approach generalizes beyond the vision domain, we further evaluate it on a representative NLP benchmark. Specifically, we fine-tune DistilBERT base-uncased, a 6-layer transformer distilled from BERT, on the AG News dataset [27]. We set the LoRA rank to $r=8$, the sparsification ratio to $O_{k}=0.5$, and maintain a consistent aggregation schedule to isolate the effect of our method. As shown in Fig. 7, our method consistently outperforms baseline schemes on this NLP task, demonstrating that the proposed technique is modality-agnostic and remains effective beyond image data. Furthermore, since our mechanism operates entirely at the LoRA update level, it can be seamlessly integrated into a wide range of LoRA-based PEFT frameworks, suggesting strong potential for on-device fine-tuning in natural language applications.
VI Conclusion
In this paper, we developed a wireless federated LoRA fine-tuning framework for communication-efficient on-device adaptation of pre-trained models. Our analysis showed that structural LoRA rank selection and the covariance induced by separate aggregation of LoRA factors play important roles in convergence. Furthermore, we proposed SOFT for efficient parameter selection and TSFA for dynamically optimizing sparsification and bandwidth. By leveraging Lyapunov optimization, TSFA ensures long-term latency stability and resource efficiency. Experiments show that the proposed framework, enhanced with SOFT and TSFA, achieves performance on par with existing methods while significantly reducing communication overhead.
Appendix A Proof of Lemma 1
Before proceeding with the main proof, we first establish essential inequality. Using Arithmetic and Geometric Mean inequality (AM-GM inequality), i.e. $\gamma a+\frac{1}{\gamma}b\geq 2\sqrt{ab}$ for $\gamma>0$, $\|x+y\|_{F}^{2}$ can be upper-bounded as:
$\|x+y\|_{F}^{2}$ $=\|x\|_{F}^{2}+\|y\|_{F}^{2}+2x^{\textsf{T}}y$ (42)
$\leq\|x\|_{F}^{2}+\|y\|_{F}^{2}+\gamma\|x\|_{F}^{2}+\frac{1}{\gamma}\|y\|_{F}^{2}$ (43)
$=(1+\gamma)\|x\|_{F}^{2}+\left(1+\frac{1}{\gamma}\right)\|y\|_{F}^{2}.$ (44)
Now, using the Assumption 1 and (14),
$\mathbb{E}\left[\left\|\tilde{m}_{k}^{t+1}\right\|_{F}^{2}\right]$ $\leq(1-O_{k})\mathbb{E}\left[\|\tilde{m}_{k}^{t}+\Delta\tilde{\theta}_{k}^{t}\|_{F}^{2}\right]$ (45)
$\leq(1-O_{k})(1+\gamma)\mathbb{E}[\|\tilde{m}_{k}^{t}\|_{F}^{2}]$
$\quad+(1-O_{k})\left(1+\frac{1}{\gamma}\right)\|\Delta\tilde{\theta}_{k}^{t}\|_{F}^{2},$ (46)
where the second inequality just follows (44). Finally, using (46) recursively and define $(1-O_{k})\left(1+\frac{1}{\gamma}\right)=\Gamma$,
$\mathbb{E}[\|m_{k}^{t+1}\|_{F}^{2}]$
$\leq\Gamma\sum\nolimits_{i=0}^{t}\left[(1-O_{k})(1+\gamma)\right]^{t-i}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ (47)
$\leq\Gamma\sum\nolimits_{j=0}^{t}\left[(1-O_{k})(1+\gamma)\right]^{j}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ (48)
$\leq\Gamma\sum\nolimits_{j=0}^{\infty}\left[(1-O_{k})(1+\gamma)\right]^{j}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ (49)
$=\frac{2(1-O_{k})(2-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ (50)
$\leq\frac{4(1-O_{k})}{O_{k}^{2}}\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2},$ (51)
where (48) is by $\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}\leq\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ and (50) is satisfied by $\gamma=\frac{O_{k}}{2(1-O_{k})}$, which ensures $(1+\gamma)(1-O_{k})<1$ for $0<O_{k}<1$.
Appendix B Proof of Theorem 1
Let $\theta_{o}$ be an original model. Then the update of the original and LoRA model at the $t+1$-th global iteration can be described as:
$\theta_{o}^{t+1}$ $=\sum\nolimits_{k\in\mathcal{N}}p_{k}\theta_{o,k}^{t+1}$
$=\theta_{o}^{t}-\sum\nolimits_{k\in\mathcal{N}}p_{k}\eta\sum_{e=0}^{E-1}\nabla F_{o,k}(\theta_{o,k}^{t,e}),$ (52)
$\theta_{r}^{t+1}$ $=\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{r,k}^{t+1}+p_{k}m_{k}^{t}+\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1})$
$=\theta_{r}^{t}-\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\eta\sum_{e=0}^{E-1}\nabla F_{r,k}(\theta_{r,k}^{t,e})+p_{k}m_{k}^{t}$
$\quad\quad+\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1}),$ (53)
where $m_{k}^{t}=m_{B}^{t}m_{A}^{t}$.
Using $S$-smoothness,
$\mathbb{E}[F_{o}(\theta_{o}^{t+1})-F_{o}(\theta_{o}^{t})]$ $\leq\underbrace{\mathbb{E}[\nabla F_{o}(\theta_{o}^{t})^{\textsf{T}}(\theta_{o}^{t+1}-\theta_{o}^{t})]}_{(a)}$
$+\underbrace{\frac{S}{2}\mathbb{E}[\|\theta_{o}^{t+1}-\theta_{o}^{t}\|_{F}^{2}]}_{(b)}.$ (54)
For $(a)$, using (52), we can obtain
$(a)$ $=-\eta\sum_{e=0}^{E-1}\mathbb{E}\left[\nabla F_{o}(\theta_{o}^{t})^{\textsf{T}}\sum_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o}^{t,e})\right]$ (55)
$=-\frac{\eta}{2}\underbrace{\sum\nolimits_{e=0}^{E-1}\mathbb{E}\left[\left\|\nabla F_{o}(\theta_{o}^{t})\right\|_{F}^{2}\right]}_{(a1)}$
$-\frac{\eta}{2}\underbrace{\sum\nolimits_{e=0}^{E-1}\mathbb{E}\left[\left\|\sum\nolimits_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o,k}^{t,e})\right\|_{F}^{2}\right]}_{(a2)}$
$+\frac{\eta}{2}\underbrace{\sum_{e=0}^{E-1}\mathbb{E}\left[\left\|\nabla F_{o}(\theta_{o}^{t})-\sum_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o,k}^{t,e})\right\|_{F}^{2}\right]}_{(a3)},$ (56)
where the second equality is due to the property $\mathbf{a}^{\textsf{T}}\mathbf{b}=\frac{1}{2}(-\|\mathbf{a}\|_{2}^{2}-\|\mathbf{b}\|_{2}^{2}+\|\mathbf{a-b}\|_{2}^{2})$. Next, we use the following lemma:
Lemma 3.
Under Assumptions 2, 3, and 6, the inequality below is satisfied:
$-\sum_{e=0}^{E-1}\mathbb{E}\left[\left\|\sum_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o,k}^{t,e})\right\|_{F}^{2}\right]$
$\leq-\mathbb{E}[\|\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}]+2S^{2}H(r_{\max}-r)W^{2}+4S^{2}\phi rW^{2}$
$\;\;+\frac{8(N-K^{t})}{K^{t}(N-1)}\eta^{2}S^{2}E^{2}G^{2}+\frac{4N(1-O^{t})^{2}}{K^{t}{O^{t}}^{4}}rS^{2}W^{4}.$ (57)
Proof:
See Appendix C. ∎
Using Lemma 3, we can derive an upper-bound on $(a2)$. Next, using the $S$-smoothness, Jensen’s inequality and the properties $\theta_{o}^{t}=\theta_{o,k}^{t,0}$ and $\sum_{k\in\mathcal{N}}p_{k}=1$, we can obtain an upper-bound on $(a3)$ as:
$(a3)$ $\leq S^{2}\sum_{k\in\mathcal{N}}p_{k}\sum_{e=0}^{E-1}\mathbb{E}\left[\left\|\theta_{o,k}^{t,0}-\theta_{o,k}^{t,e}\right\|_{F}^{2}\right]$ (58)
$=S^{2}\sum_{k\in\mathcal{N}}p_{k}\sum_{e=0}^{E-1}\mathbb{E}\left[\left\|\sum_{i=0}^{e-1}\eta\nabla F_{o,k}(\theta_{o,k}^{t,e})\right\|_{F}^{2}\right]$ (59)
$\leq S^{2}\eta^{2}\sum_{k\in\mathcal{N}}p_{k}\sum_{e=0}^{E-1}e^{2}G_{k}^{2}$ (60)
$=\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}\sum_{k\in\mathcal{N}}p_{k}G_{k}^{2}$ (61)
$\leq\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}G^{2},$ (62)
where the first equality comes from (52); the first inequality is due to Jensen’s inequality and Assumption 6; the last inequality is by $G^{2}=\max_{k}G_{k}^{2}$ and $\sum_{k\in\mathcal{N}}p_{k}=1$.
Similarly, using (53) and Jensen’s inequality, $(b)$ in (B) can be upper-bounded as:
$(b)$ $\leq\eta\mathbb{E}\left[\sum_{k\in\mathcal{N}}p_{k}\sum_{e=0}^{E-1}\left\|\nabla F_{o,k}^{t,e}(\theta_{o}^{t,e})\right\|_{F}^{2}\right]$ (63)
$\leq\eta^{2}E^{2}\sum_{k\in\mathcal{N}}p_{k}G_{k}^{2},$ (64)
$\leq\eta^{2}E^{2}G^{2},$ (65)
where the second inequality comes from Assumption 6. The last inequality is due to $G^{2}=\max_{k}{G_{k}^{2}}$ and $\sum_{k\in\mathcal{N}}p_{k}=1$.
Finally, applying the above inequalities to (B) and using $(a1)>0$ always holds, we can obtain
$\mathbb{E}$ $[F_{o}(\theta_{o}^{t+1})-F_{o}(\theta_{o}^{t})]\leq-\frac{\eta}{2}\mathbb{E}[\|\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}]+\frac{S}{2}\eta^{2}E^{2}G^{2}$
$+\eta(r_{\max}-r)HS^{2}W^{2}+2\eta S^{2}\phi rW^{2}$
$+\frac{4\eta^{3}(N-K^{t})}{K^{t}(N-1)}S^{2}E^{2}G^{2}+\frac{2\eta N}{K^{t}(O^{t})^{4}}(1-O^{t})^{2}rS^{2}W^{4}$
$+\frac{\eta^{3}E(E-1)(2E-1)}{12}S^{2}G^{2}.$ (66)
Averaging the above inequality over iteration from 0 to T1T-1 and using the property that $\theta_{o}^{0}=\theta_{r}^{0}$ since all LoRA modules are initialized with 0 at the initial stage, we can finally obtain the upper bound in Theorem 1.
Appendix C Proof of Lemma 3
First, we choose only $e=0$ from the summation of positive norm values, then we obtain
$\sum_{e=0}^{E-1}\mathbb{E}\left[\left\|\sum_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o,k}^{t,e})\right\|_{F}^{2}\right]$ (67)
$\geq\mathbb{E}\left[\left\|\sum_{k\in\mathcal{N}}p_{k}\nabla F_{o,k}(\theta_{o,k}^{t,0})\right\|_{F}^{2}\right]=\mathbb{E}[\|\nabla F_{o}(\theta_{o}^{t})\|_{F}^{2}],$ (68)
where the equality is by the initialization of the local models, $\theta_{o,k}^{t,0}=\theta_{o}^{t}$ for every client $k$. Next, we obtain $\nabla F_{r}(\theta_{r}^{t})$ by projecting $\nabla F_{o}(\theta_{o}^{t})$ onto the subspace spanned by the top $r$ singular vectors of $\theta_{o}^{t}$, thereby discarding the gradient components corresponding to the lower singular values. Since $\theta_{r}^{t}$ lies entirely within this subspace, evaluating the gradient of $F_{o}$ at $\theta_{r}^{t}$ is equivalent to evaluating the gradient of the restricted function $F_{r}$. Using a simple property, $\|\nabla F_{o}(\theta_{o}^{t})\|_{F}^{2}\geq\|\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}$, we have
$-\mathbb{E}[\|\nabla$ $F_{o}(\theta_{o}^{t})\|_{F}^{2}]\leq-\mathbb{E}[\|\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}]$
$+\underbrace{\mathbb{E}[\|\nabla F_{o}(\theta_{o}^{t})-\nabla F_{r}(\theta_{r}^{t})\|_{F}^{2}]}_{(c)}.$ (69)
Then, from $S$-smoothness, $(c)$ is upper bounded as
$(c)$ $\leq S^{2}\mathbb{E}[\|\theta_{o}^{t}-\theta_{r}^{t}\|_{F}^{2}]$ (70)
$\leq 2S^{2}\underbrace{\mathbb{E}\left[\left\|\theta_{o}^{t}-\hat{\theta}_{r}^{t}\right\|_{F}^{2}\right]}_{(c1)}+2S^{2}\underbrace{\mathbb{E}\left[\left\|\hat{\theta}_{r}^{t}-\theta_{r}^{t}\right\|_{F}^{2}\right]}_{(c2)}$ (71)
where the second inequality is due to $\|a-b\|_{F}^{2}\leq 2\|a\|_{F}^{2}+2\|b\|_{F}^{2}$ and $\hat{\theta}_{r}^{t}=\sum_{k\in\mathcal{N}}p_{k}\theta_{B,k}^{t}\theta_{A,k}^{t}$. For $(c1)$,
$(c1)$ $\leq\mathbb{E}\left[\sum_{k\in\mathcal{N}}p_{k}\left\|\theta_{o,k}^{t+1}-\hat{\theta}_{r,k}^{t+1}\right\|_{F}^{2}\right]$ (72)
$\leq H(r_{\max}-r)W^{2},$ (73)
where the first inequality comes from Jensen’s inequality and the second inequality comes from Assumption 2 and Assumption 6. Next, plugging (53), $(c2)$ can be expressed as:
$(c2)$ $=\underbrace{\mathbb{E}\left[\left\|\sum_{k\in\mathcal{N}}p_{k}\theta_{r,k}^{t}-\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{r,k}^{t}\right\|_{F}^{2}\right]}_{(c21)}$
$+\underbrace{\mathbb{E}\left[\left\|\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}m_{k}^{t}-\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1})\right\|_{F}^{2}\right]}_{(c22)}$
$+2\bigg\langle\sum_{k\in\mathcal{N}}p_{k}\theta_{r,k}^{t}-\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{r,k}^{t},$
$\quad\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}m_{k}^{t}-\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1})\bigg\rangle,$ (74)
where the last term is zero due to unbiasedness.
We follow the same steps as in Lemma 5 of [25] to derive the upper-bound $(c21)$, which utilize the probabilistic sampling properties, model unbiasedness and $\mathbb{E}[\|x-\mathbb{E}[x]\|_{F}^{2}]\leq\mathbb{E}[\|x\|_{F}^{2}]$.
$(c21)\leq\frac{4(N-K^{t})}{K^{t}(N-1)}\eta^{2}E^{2}G^{2},$ (75)
where $G^{2}=\max_{k}G_{k}^{2}$. Next, we use a simple inequality of $\|\mathbf{a}\!-\!\mathbf{b}\|_{F}^{2}\!\leq\!2\|\mathbf{a}\|_{F}^{2}\!+\!2\|\mathbf{b}\|_{F}^{2}$, Then, $(c22)$ can be bounded as
$(c22)$ $\leq 2\mathbb{E}\left[\left\|\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}m_{k}^{t}\right\|_{F}^{2}\right]$
$+2\mathbb{E}\left[\left\|\textit{Cov}(\theta_{B}^{t+1},\theta_{A}^{t+1})\right\|_{F}^{2}\right]$ (76)
$\leq\frac{2N}{K^{t}}\mathbb{E}\left[\frac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\|m_{k}^{t}\|_{F}^{2}\right]+2\phi rW^{2}$ (77)
$\leq\frac{4N(1-O^{t})^{2}}{K^{t}{(O^{t})}^{4}}rW^{4}+2\phi rW^{2}$ (78)
where the second inequality comes from Jensen’s inequality and Assumption 3; the last inequality comes from Corollary 1. Note that for the unbiasedness update, we set the sparsification ratio to be equal across the clients.